### v1

# ArtA: Automating Design Space Exploration of Spin Qubit Architectures

Abstract
In the fast-paced field of quantum computing, identifying the architectural characteristics that will enable quantum processors to achieve high performance across a diverse range of quantum algorithms continues to pose a significant challenge. Given the extensive and costly nature of experimentally testing different designs, this paper introduces the first Design Space Exploration (DSE) for quantum-dot spin-qubit architectures. Utilizing the upgraded SpinQ compilation framework, this study explores a substantial design space comprising 29,312 spin-qubit-based architectures and applies an innovative optimization tool, ArtA (Artificial Architect), to speed up the design space traversal. ArtA can leverage seventeen optimization method configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute force approach while maintaining the same result quality. After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests universal architectural features that perform optimally across all examined circuits, emphasizing the importance of maximizing quantum gate parallelization at the expense of more crosstalk interference.
I Introduction
As the field of quantum computing rapidly advances, different qubit technologies exhibit unique hardware and performance characteristics. It still remains to be seen which one (e.g., superconducting, trapped ions, quantum dots, photonics, defect-based on nitrogen-vacancy diamond centers) will succeed in scaling up quantum computing systems with high-quality qubits [55, 11]. Among them, spin qubits in quantum dots emerge as a compelling avenue for achieving scalability for practical quantum computation [11]. To this day, spin qubits are at an early stage in their development, with Intel’s Tunnel Falls chip currently boasting the highest count of twelve spin qubits [32]. Despite this, their inherent scalability advantages suggest that a robust two-dimensional design [63, 36, 7] could be scaled up relatively easily once fabrication techniques and quality improve up to a certain level. Designing a chip, however, involves multiple architectural design choices whose impact in the future can only be fully assessed through experimental studies post-scaling the technology. Exploring all these design possibilities simultaneously can be expensive, time-intensive, and impractical. Therefore, simulating how different design choices can affect performance in a range of quantum applications is crucial to assist the development. This process facilitates highlighting architectural insights that can guide the technology forward and will allow quantum researchers to make more informed decisions, streamline efforts, and hasten the development of these devices.
Recognizing the intricate challenges of designing quantum processor architectures, this study initiates the first Design Space Exploration (DSE) for spin-qubit architectures, both from a current technological perspective and a future one. Therefore, in this work, we have identified a wide range of representative architectural features and abstracted them into usable input variables, resulting in $29,312$ different architectures. To facilitate this exploration, we have enhanced the compilation capabilities of SpinQ compilation framework [47] to handle all these input variables and updated the definition of the Estimated Success Probability (ESP) metric to include not only operational errors but also crosstalk and decoherence errors. These transformations and many more establish SpinQ as the first compilation and DSE framework for spin qubit architectures.
DSE processes [50, 39], which rely on a brute force approach to traverse and subsequently analyze the design space, will be impractical time-wise, especially for such large spaces. Our approach, however, employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods designed to automate the DSE process of SpinQ, thus taking less time than the brute force approach with the same quality of results. The abilities of ArtA are (a) to suggest which of the seventeen optimization configurations can find the architecture with the desired ESP the fastest compared to the brute-force approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per quantum benchmark circuit.
Our results demonstrate ArtA’s ability to compare all optimization techniques for each quantum circuit and obtain a solution up to 99.1% faster, on average, than brute-forcing. Then, equipped with these insights, we move on to conduct a DSE analysis of $29,312$ spin qubit architectures and provide valuable insights into best design practices. Firstly, we highlight the critical need to maximize the parallelization of quantum gates rather than minimizing crosstalk between qubits. Secondly, we find that the communication method with shuttle operations is better in large-scale circuits than SWAPs, but for single-qubit gates, pulse-based methods rotations are preferred over shuttle-based. Finally, our universal architecture for all used quantum circuits suggests that the likelihood of success increases when prioritizing the parallelization of single-qubit gates instead of two-qubit gates.
The main contributions of this paper:
  1. 1.
    The design space definition comprising of abstracted characteristics of current and, potentially, future spin qubit architectures.
  2. 2.
    The new SpinQ, the first compiler and DSE framework for spin qubit architectures. In this version, we have updated the ESP formula to include operational, crosstalk, and decoherence-induced errors.
  3. 3.
    ArtA, the first tool consisting of seventeen optimization method configurations automating the DSE process of spin qubit architectures.
  4. 4.
    Evaluation of best-matching optimization configurations per quantum circuit.
  5. 5.
    Evaluation of best-matching architectural characteristics per quantum circuit and universally (i.e., for all used circuits) in terms of ESP.
The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the resulting performance of the final architecture. Then, in Section III, we motivate the need for an automated Design Space Exploration analysis for spin qubit devices and formulate four research questions for this work. After that, in Section IV, we define the design space consisting of various input architectural variables and establish the new DSE functionality of SpinQ. Another upgrade in SpinQ is the enhanced Estimated Success Probability metric, described in Section V, which is used as the Figure of Merit for our exploration. In Section VI, we introduce ArtA (Artificial Architect) and two new metrics to evaluate its performance. In the results Section VII, we first evaluate ArtA’s performance across all used quantum circuits and determine the best optimization method configuration. Then, we conduct a detailed analysis of the best architectural designs for each quantum circuit and form valuable insights for a universal architecture. We conclude our work and provide ideas for future directions in Section IX.
II Challenges for scaling up spin qubit devices
Spin-qubit technologies are distinguished by their unique physical features, which position them as a highly scalable solution for quantum computing. The advantages of spin-qubits include a significantly smaller size — up to a thousand times less than other qubit technologies — combined with decades of semiconductor manufacturing expertise, long coherence times coupled with short gate durations, and high temperatures [71, 10, 30, 11, 74, 38, 63, 66, 72, 68]. At the core is the quantum dot, which contains a trapped electron(s) or hole(s) to form a physical qubit [26]. Spin qubits are manipulated electromagnetically using multiple precision-engineered gate electrodes that facilitate either single- or two-qubit operations through exact timing of pulse sequences across various quantum dot configurations. Recent studies have expanded these systems into one-dimensional and two-dimensional arrays [30, 6], exploring different structures and material combinations.
Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and availability. Major technological hurdles are related to the so-called interconnect bottleneck [63], and various fabrication challenges towards scaling up [9, 5, 18, 17]. On the upside, there have been significant efforts [36, 8, 63, 31, 21, 45, 48, 65] to tackle these challenges.
However, solving them can not guarantee successful quantum algorithm executions. This is because the quality and quantity of qubits are not the only factors determining a high-performing quantum processor. The architectural constraints qubits need to comply with are equally important in how they are operated. For instance, the benefits of qubits with excellent operational fidelity can easily be outweighed by low qubit connectivity and limited natively supported quantum gates. Similarly, high-quality qubits with low crosstalk interference are not enough when they can not be addressed in parallel. These, and many more, are interlinked architectural trade-offs that affect the actual performance of the quantum processors.
During the Noisy Intermediate-Scale Quantum (NISQ) era [51], predicting which architectural features and trade-offs will facilitate successful quantum circuit executions while maintaining reasonable hardware requirements remains challenging. To date, there has not been a systematic study exploring spin-qubit architectures through this prism of providing concrete guidelines for future development. Therefore, in this study, we exploit such an opportunity and alleviate the need for time-consuming, expensive, and technology-dependent experimental studies.
III Problem Statement
The need to apply design space exploration (DSE) techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2]. These techniques have been successfully applied across different levels of the quantum computing stack. For instance, DSE has been used to explore ion-trap quantum processors [41] and evaluate compilation techniques for superconducting qubits [54]. Such techniques can abstractly convert architectural characteristics into design variables and, in this way, quantize the design space, expressing current and future design possibilities that otherwise would be impossible or too time-consuming to physically realize. Starting from there [50, 39], the steps for a proper DSE: i) Describe the problem in terms of input variables or parameters (design choices), ii) select the performance metrics, iii) choose a global cost function as a Figure of Merit that combines different performance metrics, and iv) find models (behavioral, analytical, experimental data) that connect input parameters with metrics (or directly to the Figure of Merit). By using these DSE techniques, performance trends can be observed projecting current or future multidimensional design spaces, and points of optimality can be identified. From such an approach, it can be determined which architectural design choices (i.e., combinations of input parameter values) could provide better performance when benchmarked from small to large-scale quantum circuits.
Although this methodology is well established and understood in many disciplines, it can be challenging to implement in such early stages for spin-qubit technologies. One of the most difficult aspects is the development of a simulation framework incorporating a range of representative architectural variables adopting a large-scale perspective. Firstly, it is difficult to predict which architectural features will be relevant for many generations of devices to come, and secondly, the framework itself has to be flexible to incorporate new ones easily.
In practice, this can be very time-consuming, and thus, typically, the design space is kept at a reasonable and, usually, small scope under specific technological assumptions. Considering the rapid progress of spin-qubits, the potential design space can be enlarged quickly with new considerations, rendering any previous exploration obsolete. Hypothesizing such a conventional DSE framework exists, it can be tedious to brute force thousands of designs and extensively analyze results. Hence, the relevancy of a framework automating such a DSE process, which can simplify performance evaluation of different spin-qubit architectures, is prominent.
To capture the above problem statement, we ask the following four research questions:
  1. 1.
    What characteristics of spin-qubit architectures are representative, and how can they be incorporated in a DSE framework? Answered in Section IV
  2. 2.
    How can a DSE framework for investigating preferred spin-qubit architectural characteristics be constructed? Answered in Section IV
  3. 3.
    Which optimization methods are best suited for automating the DSE? Answered in Section VII.1
  4. 4.
    Which architectural characteristics in the selected design space are key for building high-performance spin-qubit devices for specifically given quantum circuits? Answered in Section VII.2
IV SpinQ and the design space
Table 1: Summary of values of the considered architecture variables.
To answer the first two research questions, we conducted an analysis of relevant spin-qubit architectural properties with a long-term view and distilled them into numerous variables. These properties were selected to be representative of spin qubit devices as well as to create trade-offs between them. The latter will prevent maximizing one without causing a performance penalty through another (refer to Section V). The defined design space is summarized in Table 1, including their possible values. It should be noted that the DSE analysis in this work uses either discrete or categorical variables [42], as we aim to provide performance trends and rankings between optimization algorithms and architectures rather than absolute comparisons. An architecture is then characterized by one (valid) combination of variable values. Below, we describe the considered design space.
IV.1 Operational gate constraints
The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types (i.e., executed simultaneously) and (b) how many of them can be parallelized at the same time step. To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk.
  • a)
    Combinations of gate types that can be parallelized:
    (xy_z, xy_tqg, z_tqg, xy_z_tqg = 0, 1)
    As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel. For example, xy_z defines if X/Y can be executed simultaneously with Z gates.
  • b)
    Constraints affecting the parallelization per gate type expressed in percentage:
    (xyD, zD, tqgD, sD = -1, 1, 25%,50%, 75%, 100% )
    These variables indicate how many gates, for each type, can be parallelized, expressed as a percentage relation to the total number of qubits. For example, xyD = 25 in a 40-qubit architecture means that at most ten X or Y gates can exist in a single cycle. Note that the notion of a cycle refers to the basic unit of time representing one step in a sequence of gates of a quantum circuit, and each step may contain multiple gates. The zD and sD variables indicate the same for Z rotations and shuttle gates, respectively. The tqgD variable indicates the parallelization amount for two-qubit gates. As these always involve two qubits, they are expressed as a percentage of half of the total number of qubits. A value of $-1$ means that the parallelization degree is not user-defined but dictated by a specific single-qubit gate implementation explained next.
  • c)
    Single-qubit gate implementation:
    (single_qubit_impl = 0 $\rightarrow$ Sequential, 1 $\rightarrow$ Local, 2 $\rightarrow$ Global, 3 $\rightarrow$ Semi-global)
    This variable indicates in what manner single-qubit gates are carried out. A Global single-qubit gate implementation means the same rotation axis and angle are applied to all qubits. For qubits that do not participate in a cycle, locally applied disable instructions prevent their rotation. Then, the Semi-Global implementation is considered with the same rotation scheme as in [28] and [47], which is essentially equivalent to the implementations of the crossbar architecture proposed of [36]. In this implementation, qubit rotations are implemented semi-globally, meaning that either all qubits in odd or even column parities can be rotated at a time with the same axis and angle, and unwanted qubit rotations have to be reversed afterward by additional instructions [47]. A Local single-qubit gate implementation allows for any parallelized combination of gates at any angle to be executed in a cycle. Finally, a Sequential single-qubit gate implementation does not allow any gate parallelization, meaning that every cycle will be occupied by one gate only. It should be noted that depending on the single_qubit_impl value, only certain combinations of categories (a) and (b) are possible. For instance, in the Global and Semi-Global implementations, the xyD variable is not applicable as these implementations predetermine the upper limit of single-qubit gates allowed during a cycle.
IV.2 Implementation of Z rotations:
(z_rot_impl = 0, 1)
This boolean variable also stipulates the implementation of Z rotation gates on a hardware level. The first option is a high-fidelity shuttle-based Z rotation, which is achieved with two time-sensitive qubit shuttles to and from a neighboring column [47, 28, 40, 36]. This single-qubit manipulation belongs to the so-called “hopping spins” type [67], and it is possible to achieve an arbitrary rotation axis, but in this work, we assume the aforementioned implementation proposal for simplicity. The second option is a regular pulse-based Z rotation, where Z gates are implemented similarly to X/Y gates by applying pulses. With the first choice being a unique characteristic of spin-qubits, we are interested in exploring which one of the two provides better performance across many quantum circuits. Similarly to the variables mentioned earlier, only certain combinations are possible depending on single_qubit_impl values. For example, with a Semi-Global single-qubit gate implementation and a pulse-based Z rotation gate, the zD variable is not applicable for the same reason the xyD variable is not applicable, as we explained before in category (b).
IV.3 Connectivity of the device:
(degree = 4, 6, 8)
This variable stipulates the considered coupling graphs by addressing them through their average node degree. A coupling graph is a graph of possible direct interactions (edges) between qubits (nodes). The first coupling graph we used is a square grid (degree = 4). The second is identical to the first but adds one diagonal connection in each square (i.e., each qubit is coupled with six neighboring qubits and with three at the corners, degree = 6). Finally, the third adds the second diagonal in each square (i.e., each qubit is coupled with eight neighboring qubits and with three at the corners, degree = 8). Figure 1, presents the coupling graphs considered with color-coded edges for each of the three types. We have now reviewed all the variables related purely to hardware. Next, we will consider two variables associated also with the compilation process. More specifically, they change the routing process based on the hardware communication constraints.
Figure 1: The different coupling graphs considered in the design space. The coupling graph consists of the black lines and edges when degree = 4, black and green lines for degree = 6, and all lines when degree = 8.
IV.4 Routing methods:
(router = 0 $\rightarrow$ Shuttle-based SWAP, 1 $\rightarrow$ beSnake)
There are two routing algorithm choices. The first one, introduced in [47, 40], is a shuttle-based SWAP routing algorithm. In particular, this algorithm was tailored around the unique constraints of the crossbar architecture [36], which necessitated the maintenance of the checkerboard physical qubit pattern to achieve a time-efficient compilation process. One of the advantages of this algorithm is, in fact, the maintenance of the checkerboard pattern, which in turn keeps the crosstalk effect low and enables the Semi-Global rotations scheme exclusively. The second routing option is called beSnake [46], and it is capable of freely shuttling qubits around any topology and in any direction, handling complex routing scenarios involving parallelized gates.
IV.5 SWAP replacement:
(SWAP_opt = 0, 1)
This binary variable concerns an optional functionality of the beSnake routing algorithm, which can replace a sequence of shuttles with a SWAP operation under certain conditions [46]. These conditions are satisfied when the accumulated fidelity of a parallelized shuttle sequence exceeds the fidelity of only executing a SWAP on that particular location. This variable is useful for creating insights for architectures that support SWAPs compared to others that do not support them.
IV.6 Design interdependencies and verification
Figure 2: Tree of valid architectural configurations that SpinQ can handle
All these variable values act as input into the upgraded SpinQ housing its new configurable compiler, making it the first compilation and DSE framework for spin-qubit architectures. After this upgrade, we are able to provide SpinQ with multiple combinations of variable values, compile each input circuit iteratively for all valid architectures, and calculate and store the same performance metrics used in [47]. These are gate overhead, circuit depth overhead, and Estimated Success Probability (ESP) of the compiled quantum circuit(s).
As mentioned, not all combinations of variable assignments are valid, as there are interdependencies between them; hence, SpinQ automatically filters the design space to allow only valid combinations. If all combinations were allowed, presented in Table 1, there would be a total of $1,105,920$ architectures, but the valid ones used in this work amount to $29,312$. Figure 2 shows the interdependences tree that SpinQ uses to filter valid architectures.
Another significant component of SpinQ is the compiler verification tool [47]. This tool is crucial, given the absence of real devices for testing. The verification functions have been enhanced to handle all valid architectures via our two-step method, which thoroughly scrutinizes circuits at each compilation stage [47].
V Figure of Merit
To estimate the performance of any quantum processor, a prevalent method when the device is not physically available is through the Estimated Success Probability (ESP) [53, 43] extracted from the compiled circuit. This approach is by far the most efficient compared to Schrodinger-Feynman, tensor networks, or noise model simulations as all become exponentially slower with larger circuits and/or higher qubit connections on the device[47]. The choice of the figure of merit is a crucial step to ensure fair results, as every design variable needs to be directly linked to the resulting value, along with their trade-offs. This will prevent ArtA (introduced later in Section VI) from maximizing some variable(s) and, by doing so, unfairly favoring specific architectures. For example, it will not be fair to be able to maximize the parallelization degree of all gates without causing some negative impact due to increased crosstalk. Such a consideration also expresses more accurately design decisions in real experimental quantum processors.
In this work, we start from the ESP definition of [47] given in Equation 1 and expand it to reflect all previous considerations, given in Equation 2, with now three components in total: a gate operations term, a crosstalk term, and a decoherence term.
$ESP=\prod_{k}\prod_{i}{F_{i,k}}$ (1)
where $k$ represents the $k$th time step and $i$ the $i$th gate in the $k$th time step.
$ESP=(\prod_{k}\prod_{i}{F_{i,k}})\cdot(\prod_{k}\prod_{i,j}C_{i,j,k})\cdot(1-e^{(-t/T_{2})})^{N}$ (2)
where $k$ represents the $k$th time step, $i$ the $i$th gate in the $k$th time step and $j$ represents the $j$th gate that inflicts crosstalk with the $i$th gate. $F$ is the fidelity and $C$ the crosstalk of the corresponding gates. The last term, adapted from [28], introduces the decoherence-induced errors, which represent the probability of all qubits staying coherent during the execution of the circuit. $T_{2}$ is the decoherence time, $t$ the duration time of the circuit, and $N$ the total number of qubits. We will discuss each of the terms in detail in the next three subsections.
V.1 Gate operations term
The choice of typical operational fidelities of each gate, coined by the $F_{i,k}$ term, plays a large role in the probability of success of quantum circuits. In our analysis, we consider these fidelities to be constants within the design space rather than variables. This is because determining the optimal fidelity value for enhancing the ESP is straightforward; higher operational fidelity invariably leads to improved performance. Since we are conducting an exploration, we take a highly optimistic approach to the state-of-the-art [49, 70, 44] values for each gate type [^1] Another reason we do so is to avoid numeric underflows where there is a loss of accuracy in numerical calculations if ESP becomes smaller than the smallest positive representable value in a programming language’s floating-point arithmetic.:
$F_{\text{single-qubit}}=0.9999,F_{\text{two-qubit}}=0.9998$ (3)
V.2 Crosstalk term
Crosstalk has been sparsely studied for spin qubits. This is because the crosstalk effects when executing multiple gates in parallel in close proximity are still very complex to model accurately [61, 27]. Nonetheless, the presence of crosstalk is expected to have a large influence on quantum computer performance in the future, and current proposals are already suggesting techniques to mitigate it in simple forms [61]. For our purposes, crosstalk creates necessary trade-offs between architectural variables, and, as explained before, this is essential for a fair ESP calculation across architectures. For these reasons, we have devised a model to help us discover relative categorical differences between architectures, rather than accurate predictions of crosstalk effects. Our overall aim is not to approximate the actual performance of designs but to pronounce their differences by creating a performance order between them. We further assume that crosstalk should depend on local hardware characteristics between any two nearest-neighboring qubits being operated at the same time, and since this information is already included in the operational fidelity of each gate type, our crosstalk definition is a function of those. Below, we summarize our assumptions:
  • Nearest-neighbor crosstalk only (i.e., a pair of qubits that are directly connected by an edge in the coupling graph)
  • Crosstalk occurs for each edge connecting qubits, which are operated in parallel by different gates
  • There should be some correlation between the fidelity of gates and crosstalk effects, as they are both influenced by the same quality of fabrication
  • The crosstalk occurring between neighboring qubits on which two different quantum gates are applied should not be higher than the combined operational error of these two gates
Based on these assumptions, we calculate the crosstalk effect as follows:
$C_{i,j,k}={(\frac{2}{\frac{1}{F_{i,k}}+\frac{1}{F_{j,k}}})}^{n}$ (4)
where $n$ is the number of the total direct links (i.e., nearest-neighbors) on the topology between the $i$th and the $j$th operation. Similarly to before, $F$ corresponds to the operational fidelity.
A schematic overview of crosstalk effects is given in Figure 3. We see two two-qubit gates acting within the red ovals between Q-1 and Q-2, and Q-3 and Q-4, respectively. In this hypothetical square grid with one diagonal topology, crosstalk occurs in all three edges marked with a flash, as they directly connect the qubits of these two operations.
Figure 3: Schematic overview of cross-talk when executing two-qubit gates. Dashed black lines indicate a direct connection between two qubits, the red ovals indicate an ongoing two-qubit gate, and the blue flash icons indicate contributing crosstalk effects.
V.3 Decoherence term
The last term of Equation 2 represents the decoherence of quantum information; hence, it creates a strong trade-off between parallelization and crosstalk effects. Since $N$ and $T_{2}$ are given, we need to define the duration time of the circuit $t$:
$t=\sum_{k=1}^{M}\text{max}(D_{i,k})$ (5)
where $D_{i,k}$ is the duration of the $i$th operation in the $k$th time step. In Table 2, we present the gate duration and decoherence time used. The motivation behind this selection of durations, once again, is for purposes of architectural comparisons. We took reference for our two-qubit gate (i.e., $\sqrt{SWAP}$) from [44, 70], for the $Shuttle$ from [28] which is half the shuttle-based Z rotations [47], for the pulsed-based single-qubit gates from [49], and the $T_{2}$ from [29].
Table 2: Overview of all gate duration and the decoherence time used.
VI ArtA (Artificial Architect)
Figure 4: Schematic overview of the initialization, optimization, and output cycle of ArtA (red box) with SpinQ (blue box).
From the aforementioned transformation of the SpinQ compilation framework to a Design Space Exploration (DSE) framework, we are able to brute force through all valid combinations of the input variable values. However, as we discussed, this can be a time-consuming process, especially when the defined design space is large. Therefore, a specialized tool is needed to automate the DSE process by traversing the space in a clever and faster way. Ultimately, this will result in fewer designs being tried until desired ESP values or other metrics are reached.
To do that, we introduce ArtA, a tool built into the SpinQ framework containing multiple optimizers. Its goal is to give answers to (a) which optimization technique performs best on the task of finding an architecture that achieves the desired ESP and (b) which architecture provides the highest ESP, both per quantum circuit.
The abstracted ArtA process is depicted in Figure 4. The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Sec. IV, the optimization method(s) to traverse the space, and an (optional) termination condition. Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture (in the form of input variable values) from the available “population” of architectures – a subset of the complete design space determined by the optimization algorithm itself. This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired value of a specific metric is reached (in this work, ESP). Upon termination, the resulting architecture, its ESP, and ArtA metrics are stored. In the following sections, we will describe the TC and ArtA metrics.
In the Appendix X.2, we analytically discuss the five optimization methods used and their hyper-parameter variations, totaling seventeen different optimization configurations. Also, in the Appendix X.1, we present the seven benchmark quantum circuits and their different sizes in terms of qubit counts.
To answer the last two research questions, we will use ArtA’s (a) and (b) capabilities. In the first one, we answer the third research question by testing all the implemented optimization methods to determine which finds the highest ESP the fastest for each quantum circuit. Therefore, we will compile each circuit for all $29,312$ architectures and store their ESP. Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Sec. VI.2 evaluating ArtA’s overall performance. In the second way, we will only use the best-performing optimization method but expand our circuit selection to concretely answer the last research question.
VI.1 Termination Condition (TC)
During each optimization cycle, a specific termination condition (TC) is checked to reduce further computational time based on the way ArtA is used. Since we already obtained the ESP values for all architectures first during the third question, the highest ESP is known and can work as a TC. Additionally, we observed in practice that most runs reach the highest ESP sooner than 40 minutes; hence, this can be another TC. More specifically, our analysis revealed that over 95% of processes are completed earlier, while a minority extends significantly longer. Given that our evaluation focuses on runtime, omitting these prolonged instances won’t affect the overall assessment, as their inclusion would lead to unfavorable evaluations regardless.
Therefore, the TC when answering the third research question is:
  • If the highest ESP value is found.
  • If the method’s run time exceeds 40 minutes.
For the second use of ArtA, which answers the fourth research question, we investigate the returned architectural designs by using the chosen optimization method derived from answering the third question. We find in practice that the corresponding chosen optimization algorithm did not exceed trying 23.4% of the total $29,312$ architectures in the worst case before finding the one with the highest ESP.
Therefore, when answering the last research question, the TC triggers:
  • When the number of evaluated architectures exceeds 23.4% of the total $29,312$.
VI.2 Metrics Evaluating ArtA
To determine the performance of ArtA in the context of replying to the third research question, we need to define some relevant metrics. The optimization methods that will eventually determine ArtA’s performance should return a desired solution faster compared to searching with the brute-force approach. Otherwise, ArtA will be of no practical interest. We thus use the relative time-to-solution as a metric in Equation 6. The numerator includes the total runtime of both the optimization method and compilation until the desired result is reached (or the TC is triggered), and the denominator refers to the total runtime to brute force the entire design space.
[…]
When the average compilation time increases due to larger quantum circuits, the execution time of the optimization method slowly becomes less important within the total runtime. Because of that, time-to-solution is not sufficient in evaluating ArtA’s performance. Thus, a relative calls-to-solution metric, in Eq. 7, is also a strong performance predictor for ArtA, especially for larger circuits. With calls-to-solution in the numerator, we refer to the total number of times SpinQ is called to compile for each new architecture. In case the optimization method attempts the same architecture twice or more, the compilation will be skipped since the results are already stored from the first call. Finally, the denominator represents the brute force approach and is the total number of architectures in the design space, which essentially equals $29,312$.
[…]
VII Results and Evaluation
In this section, we will attempt to completely answer the third and fourth search questions. In both questions, we will focus on finding the architecture that obtains the highest ESP among all in the design space. However, before that, we briefly detail our experimental setup. Since all optimization methods are inherently probabilistic, multiple runs are required to create robust conclusions. Due to limited computational resources [^2] All time metrics were retrieved by running ArtA with Python 3.10.8 single-threaded on a 2017 MacBook Pro, using a 3.1 GHz Quad Core Intel i7 processor and 16 GB of RAM., we ran each method for each hyper-parameter combination ten times alongside the TC. Besides the mean performance, we will also report the worst performing results for reference.
VII.1 Comparison of optimization methods
Table 3: Relative time-to-solution metric (the lower, the better) for finding the architecture with the highest ESP from the design space. A minus symbolizes that there was at least one run (out of the ten) in which the top ESP value was not reached in time (TC = 40 minutes). These results are colored red, as well as runs where the time taken exceeds the brute force time (values over 100). The green cells indicate the top five configurations of optimization methods for a given quantum circuit. The mean and worst performing configurations from a sample of ten runs are shown for each quantum circuit (block of rows). QFT: Quantum Fourier Transformation, cA: Cuccaro Adder, vA: vbe Adder, QV: Quantum Volume, G: Grover’s, RC: Random Circuit, RS: Random Sampling, SA: Simulated Annealing (Top: Start Temperature, bottom: Step Size), BO: Bayesian Optimization (Top: Acquisition function, bottom: Kernel), GA: Genetic Algorithm (Top: Population Size, bottom: Mutation), ACO: Ant Colony Optimization (Top: Population Size, bottom: Exploitation Rate).
In this first part of our results, we will address the third research question stated in Section III, which is about finding the best performant optimization method according to our ArtA metrics defined in Section VI.2. The results of the relative time-to-solution metric for finding the architecture with the highest ESP are given in Table 3. This table also provides insights into matching different method configurations (described in Appendix X.2) with specific quantum circuits. As explained previously, the denominator of Eq. 6 is obtained beforehand by measuring the time taken by the brute force approach. We observe overall that for each quantum circuit, on average, all optimization methods find the solution in less time than brute-forcing (values under 100). For a better interpretation of the numbers on the table, a 50 means the particular optimization configuration of ArtA takes half the time of brute forcing all architecture, and 100 means that it takes the same time.
Multiple optimization methods greatly outperform the random sampling (RS column), as expected. The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% on average of the total brute force time to find the architecture with the highest ESP.
We calculated based on Table 3 that, on average, for all optimization methods, ArtA was able to decrease the exploration time by 80.6%. Overall, different optimization methods clearly show varied performance across their respective columns. However, the performance differences among various hyper-parameter settings within a single optimization method are less distinct. In a performance order, we can observe GA being first, followed by BO, ACO, and, lastly, SA. More specifically, GA, with a population size of 50 and a mutation probability of 0.15, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, MA] coming close at 12.1%. Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results. To further support our findings, next, we will assess the relative calls-to-solution metric.
The results of the relative calls-to-solution metric for reaching the highest ESP are given in Table 4. We observe overall that for each quantum circuit, all optimization methods were able to find the highest ESP architecture in fewer compiler calls than brute-forcing through all $29,312$ architectures. Once again, for a better interpretation of the numbers on the table, 50 means the particular optimization configuration of ArtA needs to go through half the architectures in the design space, and 100 means it needs to go through all the architectures.
The best optimization method only takes 1% of the total compiler calls, on average, for QV8 and QV10 with ACO [50, 0.1]. We calculated based on Table 4 that, on average, ArtA’s optimization method configurations can decrease the compiler calls by 79.9%. We can observe, once again, that BO, followed by GA, obtained the best performance compared to other optimization methods. Lastly, in 91.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, showing results consistency again.
Table 4: Relative calls-to-solution metric (the lower, the better) for finding the architecture with the highest ESP from the design space. A minus with red color cells symbolizes that there was at least one run (out of the ten) in which the top ESP value was not reached in time (TC = 40 minutes). The green cells indicate the top five configurations of optimization methods for a given quantum circuit. The mean and worst performing methods from a sample of ten runs are shown for each quantum circuit (block of rows). QFT: Quantum Fourier Transformation, cA: Cuccaro Adder, vA: vbe Adder, QV: Quantum Volume, G: Grover’s, RC: Random circuit, RS: Random Sampling, SA: Simulated Annealing (Top: Start Temperature, bottom: Step Size), BO: Bayesian Optimization (Top: Acquisition function, bottom: Kernel), GA: Genetic Algorithm (Top: Population Size, bottom: Mutation), ACO: Ant Colony Optimization (Top: Population Size, bottom: Exploitation Rate).
From the different variants of the GA and BO used, we focus on GA [100, 0.2] and BO [EI, MA], as they have the best performance and did not reach our TC in any of the ten runs. By comparing the two tables, we can observe that BO [EI, MA] needs marginally more time to evaluate fewer architectures compared to GA [100, 0.2]. Since the compilation time per architecture is the same between methods, this points to a slightly (0.2% on average) longer execution time of BO [EI, MA] compared to GA [100, 0.2]. This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Sec. X.2). However, the actual total runtime of ArtA will depend on the ratio between compilation and optimization time, which can change for large circuits. Based on their small relative time-to-solution difference, we can predict that BO [EI, MA] will outperform GA [100, 0.2] for larger circuits as the latter has, on average, 58% higher relative calls-to-solution. Therefore, we conclude that BO [EI, MA] is a better candidate for automating the DSE process for large-scale circuits, as it can traverse the design space faster.
Having said that, making a more detailed analysis for each circuit where internal circuit characteristics are correlated [47, 3] to optimization methods can improve the performance of ArtA by picking methods more accurately. This promising avenue is left for future work when more progress in circuit characterization is made.
VII.2 Architecture Analysis
Table 5: Architectures returned by ArtA with BO [EI, MA] optimization method for a range of quantum circuits. Results are color-coded, with color schemes following the categories introduced in Section IV.
We move on to address the last research question after having selected the best-performing optimization method in the previous section. Here, we use ArtA to create architectural insights into spin-qubit devices for each circuit category and suggest a universal architecture for all executed circuits.
In Table 5, we present the values of the architectural variables, introduced in Sec. IV, that were found by ArtA having the highest ESP with the BO [EI, MA] optimization method alongside their ESP values. These circuits are organized by class and size in number of qubits within the rows of the table. Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Sec. IV. The use of color accents in the table serves to enhance visual clarity, with each corresponding to the values of each variable.
Before focusing on the architectural insights, we should comment on the low ESP values observed, especially in larger circuit sizes. Although such low values do not have any physical meaning after a certain small number, ESP still remains a reliable way to rank architectures. It’s important to understand ESP is not inherently random, regardless of how low it gets. For instance, slight variations between two architectures causing a single gate difference will consistently and reliably be reflected through their ESPs. As long as numeric underflows are not caused, ESP remains a robust Figure of Merit for our purposes.
Observing the data presented in Table 5, it becomes evident that there is a discernible structure to the values of architectural variables. Notably, many or all circuits exhibit strong preferences for certain variables. Following an implications discussion on the circuit performance of these strong architectural preferences, we will explore areas of weaker preference where there is more variability in value choices.
Strong preference
First, we observe a unanimous preference for the Z rotation implementation (z_rot_impl) in the zri column, with a pulse-based rotation being preferred. Note that previously, we have assumed in Section V.1 that a shuttle and a pulse-based Z rotation share the same single-qubit gate fidelity, but because the latter requires two shuttles, it can benefit the ESP more than the other type. Moreover, as a shuttle operation involves two quantum dots while a pulse-based operation only involves one, the shuttle-based Z rotation will have a higher chance of incurring crosstalk errors based on our crosstalk definition. Given these two reasons, BO has favored the shuttle-based version. This indicates that shuttle-based single-qubit rotations are not as sustainable as pulsed-based for large-scale circuits.
Moving on, the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6. Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (BO [EI, MA]), as we can observe in the table. Additionally, xy_z_tqg can also be conceptually equivalent to xy_tqg and z_tqg since the pulse-based Z rotation is selected. Therefore, whenever xy_z_tqg is 1 or xy_tqg and z_tqg is 1, it means the architecture(s) which allow(s) parallelization of all single- and two-qubit gates obtains the highest ESP. In that sense, one would expect xy_z_tqg, xy_tqg and z_tqg to be synchronized. However, since such exploration freedom was given in the design space, the reasons for the mismatches could be explained by the internal characteristics of the circuits themselves. For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, which has xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it is not necessary for the hardware to support parallel executions of single-qubit and two-qubit gates. This shows that the success of real quantum circuits can come from different “angles” with reduced hardware capabilities  [^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics)..
Another strong preference appears in the beSnake router, observed by the dark green 1 values in the router column. This result shows that the crosstalk implications of using the beSnake router (which gives more shuttling freedom) over the shuttle-based SWAP router (which minimizes crosstalk) are not as important as the other two components in Equation 2. Similarly, the degree variable is mostly maximized, indicating again that the additional possibility of crosstalk does not outweigh other benefits, such as having more connections to perform two-qubit gates and shuttling operations.
Lastly, there is a strong preference for a local single-qubit gate implementation (single_qubit_impl). A local implementation, on the one hand, provides the highest parallelization possibilities, but on the other, it is the most demanding to achieve experimentally on real hardware.
Based on all the above conclusions, we can summarize that the crosstalk effect is not as strong even though ArtA suggests architectures with high crosstalk factors (the highest connectivity and parallelization and the most dense routing strategy). In fact, the decoherence-induced errors increase more severely (exponentially) with time, based on Equation 4, than crosstalk, especially in circuits with high qubit and gate counts, which explains the prevailing preference for maximizing parallelization.
The same can be observed in our universal architecture recommendation in the last row, which performs best on average for all the circuits examined. Additionally, the operational fidelity, which is directly linked to the gate overhead of compilation, can also severely impact the final ESP for circuits with high gate counts, which in turn explains the need to add the least additional gates possible during compilation through the use of beSnake [46] and local single-qubit implementation. Having said that, Bernstein Vazirani, QV8, and QV10 circuits preferred the global implementation, as these circuits consist of a significant number of similar single-qubit gates being executed at the same time step.
Weak preference
Overall, we can observe that the single-qubit operations and shuttling prefer a higher degree of parallelization than the two-qubit gates. Considering that xyD and zD have the same meaning, their combined averages (excluding -1 values) from the last row of Table 5 is 72.7% and the sD to 66.7%. This compares to an average of 52.6% for the two-qubit gates. Therefore, prioritizing the simultaneous operation of single-qubit gates (and shuttling) over two-qubit gates is more important in spin-qubit architectures.
The swap_opt variable seems to show a preference for supporting SWAP gates for lower qubit counts. A possible explanation for this is that SWAP operations take a longer time than a parallelized sequence of shuttles [46]. As we explained before, the effects of having longer circuit execution times become more severe with larger circuits; hence, the extra SWAPs can outweigh the gained fidelity they initially offer. Combining this conclusion for large-scale circuits and the strong preferences for shuttling freedom (beSnake router) indicates that the shuttle operation is a key communication method for spin-qubit devices.
VIII Acknownledgement
This work is part of the research program OTP with project number 16278, which is (partly) financed by the Netherlands Organisation for Scientific Research (NWO). CGA acknowledges funding from the Spanish Ministry of Science, Innovation and Universities through the Beatriz Galindo program 2020 (BG20-00023) and from the European ERDF under grant PID2021-123627OB-C51.
IX Conclusion
In this paper, we propose a comprehensive exploration of the current and future architectural design space for spin-qubit-based quantum dots quantum processors. Our research focuses on assessing the critical architectural characteristics that could be key in ensuring high performance on such devices. To this end, we present an upgraded version of the SpinQ compilation framework for spin qubit architectures in which architectural variables can configure internal compilation passes, making it the first Design Space Exploration (DSE) framework. After defining our design space, consisting of $29,312$ architectures, we propose a multi-optimization-based tool, ArtA (Artificial Architect), to automate the DSE process. The goal of ArtA is (a) to suggest which of the seventeen optimization method configurations can find the architecture with the highest ESP the fastest compared to the brute forcing approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per circuit category. We have shown up to 99.1% improvement in computation times compared to the brute-force approach. We have also provided insights into matching different optimization methods with specific quantum circuit categories. Bayesian optimization with an Expected Improvement acquisition function and a Matérn kernel was chosen as the best optimization method due to its effectiveness in managing higher qubit counts. After that, we proceed to automate the DSE process for 44 quantum circuits of up to 258 qubits within the 29,312-architecture design space. Our findings reveal the importance of minimizing the circuit duration by maximizing parallelization whenever possible instead of selecting architectural characteristics that minimize crosstalk errors. Crosstalk is currently a significant concern in experiments, but our long-term considerations deem decoherence more crucial. Lastly, the shuttle operation is preferred over costly SWAP operations for communication purposes, especially for large-scale circuits. Contrary to that, shuttle-based single-qubit gates such as Z rotations are not preferred over pulsed-based. Furthermore, our universal architecture suggestion places higher priority on parallelizing single-qubit gates rather than two-qubit gates.
As for future improvements, a broader range of hyper-parameter combinations and circuit variations could enhance ArtA’s effectiveness. Then, with advancements in spin qubit device fabrication, new architectural variables, and values could potentially update the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA. Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world conclusions regarding developing spin qubits. This would allow for operational fidelities to become variables as well and, consequently, add a second dimension to the Figure of Merit (ESP).
Also, some of the assumptions made in this work could be fine-tuned as the field progresses. It is crucial to acknowledge the impact of these assumptions on the observations presented. For instance, we assumed that the fidelity of pulse-based Z rotations is equivalent to that of a single shuttle operation. This assumption should not be construed as either favorable or unfavorable; rather, it is a foundational premise upon which ArtA bases its architectural “recommendations”. Should empirical evidence later demonstrate that shuttle operations exhibit significantly higher fidelity compared to pulse-based Z rotations or the crosstalk model depends on more factors, it may necessitate a reevaluation of preferences.
Further analysis of architectural patterns with similar quantum circuit characteristics can reveal more trends. For instance, quantum circuits with high two-qubit gate percentage ranges can perform best when the qubit connectivity of the architecture is above a certain degree. These kinds of statements or more can be made for a plethora of circuit characteristics [3, 47] and, in doing so, create a general matching between algorithm categories and architectural designs. Once this is done, it is possible to make the reversed suggestion of which architecture will perform best for an input circuit of interest once categorized without needing to repeat the DSE process.

### v2

@@
 Utilizing the upgraded SpinQ compilation framework, this study explores a substantial design space comprising 29,312 spin-qubit-based architectures and applies an innovative optimization tool, ArtA (Artificial Architect), to speed up the design space traversal.
-ArtA can leverage seventeen optimization method configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute force approach while maintaining the same result quality.
-After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests universal architectural features that perform optimally across all examined circuits, emphasizing the importance of maximizing quantum gate parallelization at the expense of more crosstalk interference.
+ArtA can leverage 17 optimization configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute force approach while maintaining the same result quality.
+After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests specific and universal architectural features that provide optimal performance across the examined circuits.
+Our work demonstrates that the synergy between DSE methodologies and optimization algorithms can effectively be deployed to provide useful suggestions to quantum processor designers.
 I Introduction
@@
 This process facilitates highlighting architectural insights that can guide the technology forward and will allow quantum researchers to make more informed decisions, streamline efforts, and hasten the development of these devices.
-Recognizing the intricate challenges of designing quantum processor architectures, this study initiates the first Design Space Exploration (DSE) for spin-qubit architectures, both from a current technological perspective and a future one.
+Recognizing the intricate challenges of designing quantum processor architectures, this study initiates the first Design Space Exploration (DSE) for quantum dot spin-qubit architectures, both from a current technological perspective and a future one.
 Therefore, in this work, we have identified a wide range of representative architectural features and abstracted them into usable input variables, resulting in $29,312$ different architectures.
 To facilitate this exploration, we have enhanced the compilation capabilities of SpinQ compilation framework [51] to handle all these input variables and updated the definition of the Estimated Success Probability (ESP) metric to include not only operational errors but also crosstalk and decoherence errors.
-These transformations and many more establish SpinQ as the first compilation and DSE framework for spin qubit architectures.
-DSE processes [50, 39], which rely on a brute force approach to traverse and subsequently analyze the design space, will be impractical time-wise, especially for such large spaces.
+These transformations, and many more, establish SpinQ as the first compilation and DSE framework for spin qubit architectures.
+Standard DSE processes [54, 43] will be impractical time-wise, especially for such large spaces, as they rely on a brute force approach to traverse and subsequently analyze the design space.
 Our approach, however, employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods designed to automate the DSE process of SpinQ, thus taking less time than the brute force approach with the same quality of results.
@@
 Finally, our universal architecture for all used quantum circuits suggests that the likelihood of success increases when prioritizing the parallelization of single-qubit gates instead of two-qubit gates.
+With these, we showcase that combining DSE methodologies with optimization algorithms effectively assists in forming beneficial recommendations for quantum processor designers for current and future technological stages.
 The main contributions of this paper:
   1. 1.
-    The design space definition comprising of abstracted characteristics of current and, potentially, future spin qubit architectures.
+    The design space definition which comprises five main characteristics of current and, potentially, future spin qubit architectures.
   2. 2.
@@
     Evaluation of best-matching architectural characteristics per quantum circuit and universally (i.e., for all used circuits) in terms of ESP.
-The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the resulting performance of the final architecture.
+The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the performance of the resulting architecture.
 Then, in Section III, we motivate the need for an automated Design Space Exploration analysis for spin qubit devices and formulate four research questions for this work.
@@
 Spin qubits are manipulated electromagnetically using multiple precision-engineered gate electrodes that facilitate either single- or two-qubit operations through exact timing of pulse sequences across various quantum dot configurations.
-Recent studies have expanded these systems into one-dimensional and two-dimensional arrays [30, 6], exploring different structures and material combinations.
-Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and availability.
+Studies have expanded these systems into one-dimensional and two-dimensional arrays [32, 7], exploring different structures and material combinations.
+Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and device availability.
 Major technological hurdles are related to the so-called interconnect bottleneck [68], and various fabrication challenges towards scaling up [10, 6, 19, 18].
@@
 This is because the quality and quantity of qubits are not the only factors determining a high-performing quantum processor.
-The architectural constraints qubits need to comply with are equally important in how they are operated.
+The architectural constraints qubits need to comply with in how they are operated are equally important.
 For instance, the benefits of qubits with excellent operational fidelity can easily be outweighed by low qubit connectivity and limited natively supported quantum gates.
@@
 III Problem Statement
-The need to apply design space exploration (DSE) techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2].
-These techniques have been successfully applied across different levels of the quantum computing stack.
-For instance, DSE has been used to explore ion-trap quantum processors [41] and evaluate compilation techniques for superconducting qubits [54].
+The need to apply design space exploration (DSE) techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2, 65, 5].
+These techniques or similar ones have been successfully applied across different levels of the quantum computing stack.
+For instance, they have been used to explore ion-trap quantum processors [45], evaluate compilation techniques [58] and qubit connectivities [41, 39, 20] for superconducting qubits.
 Such techniques can abstractly convert architectural characteristics into design variables and, in this way, quantize the design space, expressing current and future design possibilities that otherwise would be impossible or too time-consuming to physically realize.
@@
   3. 3.
-    Which optimization methods are best suited for automating the DSE? Answered in Section VII.1
+    Which optimization methods are best suited for automating the DSE per quantum cicruit? Answered in Section VII.1
   4. 4.
@@
 The defined design space is summarized in Table 1, including their possible values.
-It should be noted that the DSE analysis in this work uses either discrete or categorical variables [42], as we aim to provide performance trends and rankings between optimization algorithms and architectures rather than absolute comparisons.
+It should be noted that the DSE analysis in this work uses either discrete or categorical variables [46], as we aim to provide performance trends and rankings between optimization methods and architectures rather than absolute comparisons.
 An architecture is then characterized by one (valid) combination of variable values.
-Below, we describe the considered design space.
+Below, we describe the considered design space analytically.
 IV.1 Operational gate constraints
-The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types (i.e., executed simultaneously) and (b) how many of them can be parallelized at the same time step.
+The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types can be executed simultaneously and (b) how many of them can be parallelized at the same time step.
 To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk.
@@
     (xy_z, xy_tqg, z_tqg, xy_z_tqg = 0, 1)
-    As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel. For example, xy_z defines if X/Y can be executed simultaneously with Z gates.
+    As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel.
+For example, xy_z defines if X/Y gates can be executed simultaneously with Z gates.
   • b)
@@
 In this implementation, qubit rotations are implemented semi-globally, meaning that either all qubits in odd or even column parities can be rotated at a time with the same axis and angle, and unwanted qubit rotations have to be reversed afterward by additional instructions [51].
-A Local single-qubit gate implementation allows for any parallelized combination of gates at any angle to be executed in a cycle.
+A Local implementation allows for any parallelized combination of gates at any angle to be executed in a cycle.
 Finally, a Sequential single-qubit gate implementation does not allow any gate parallelization, meaning that every cycle will be occupied by one gate only.
@@
 The first option is a high-fidelity shuttle-based Z rotation, which is achieved with two time-sensitive qubit shuttles to and from a neighboring column [51, 30, 44, 38].
-This single-qubit manipulation belongs to the so-called “hopping spins” type [67], and it is possible to achieve an arbitrary rotation axis, but in this work, we assume the aforementioned implementation proposal for simplicity.
-The second option is a regular pulse-based Z rotation, where Z gates are implemented similarly to X/Y gates by applying pulses.
+This single-qubit manipulation belongs to the so-called “hopping spins” type [72], and it is possible to achieve an arbitrary rotation axis, but in this work, we only assume the aforementioned variety for simplicity.
+The second option is a regular pulse-based Z rotation, where Z gates are implemented similarly to X/Y gates by magnetic pulse interactions.
 With the first choice being a unique characteristic of spin-qubits, we are interested in exploring which one of the two provides better performance across many quantum circuits.
@@
 Next, we will consider two variables associated also with the compilation process.
-More specifically, they change the routing process based on the hardware communication constraints.
+More specifically, they change the routing process but based on the hardware communication constraints.
 Figure 1: The different coupling graphs considered in the design space.
-The coupling graph consists of the black lines and edges when degree = 4, black and green lines for degree = 6, and all lines when degree = 8.
+The coupling graph consists of the black line edges when degree = 4, black and green edges when degree = 6, and all edges when degree = 8.
 IV.4 Routing methods:
@@
 IV.6 Design interdependencies and verification
-Figure 2: Tree of valid architectural configurations that SpinQ can handle
+Figure 2: Tree of valid architectural configurations that SpinQ handles.
 All these variable values act as input into the upgraded SpinQ housing its new configurable compiler, making it the first compilation and DSE framework for spin-qubit architectures.
@@
 Our overall aim is not to approximate the actual performance of designs but to pronounce their differences by creating a performance order between them.
-We further assume that crosstalk should depend on local hardware characteristics between any two nearest-neighboring qubits being operated at the same time, and since this information is already included in the operational fidelity of each gate type, our crosstalk definition is a function of those.
+We further assume that crosstalk should depend on local hardware characteristics, and since this information is already included in the operational fidelity of each gate type, our crosstalk definition is a function of those.
 Below, we summarize our assumptions:
@@
 V.3 Decoherence term
-The last term of Equation 2 represents the decoherence of quantum information; hence, it creates a strong trade-off between parallelization and crosstalk effects.
+The last term of Equation 2 represents the decoherence of quantum information, which creates a strong trade-off between parallelization and crosstalk effects.
 Since $N$ and $T_{2}$ are given, we need to define the duration time of the circuit $t$:
@@
 VI ArtA (Artificial Architect)
-Figure 4: Schematic overview of the initialization, optimization, and output cycle of ArtA (red box) with SpinQ (blue box).
+Figure 4: Schematic overview of the initialization, optimization, and output cycle when using ArtA (red box) together with the upgraded SpinQ framework (blue box).
 From the aforementioned transformation of the SpinQ compilation framework to a Design Space Exploration (DSE) framework, we are able to brute force through all valid combinations of the input variable values.
@@
 To do that, we introduce ArtA, a tool built into the SpinQ framework containing multiple optimizers.
-Its goal is to give answers to (a) which optimization technique performs best on the task of finding an architecture that achieves the desired ESP and (b) which architecture provides the highest ESP, both per quantum circuit.
+Its goal is to give answers to (a) which optimization technique performs best on the task of finding an architecture that achieves the desired ESP and (b) which architecture provides the highest ESP, both per quantum circuit and universally.
 The abstracted ArtA process is depicted in Figure 4.
-The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Sec.
-IV, the optimization method(s) to traverse the space, and an (optional) termination condition.
+The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Section IV, the optimization method(s) to traverse the space, and an (optional) termination condition.
 Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture (in the form of input variable values) from the available “population” of architectures – a subset of the complete design space determined by the optimization algorithm itself.
-This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired value of a specific metric is reached (in this work, ESP).
+This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired value of a specific metric is reached (ESP in this work).
 Upon termination, the resulting architecture, its ESP, and ArtA metrics are stored.
-In the following sections, we will describe the TC and ArtA metrics.
+In the following sections, we will describe the TC and ArtA metrics in more detail.
 In the Appendix X.2, we analytically discuss the five optimization methods used and their hyper-parameter variations, totaling seventeen different optimization configurations.
@@
 Therefore, we will compile each circuit for all $29,312$ architectures and store their ESP.
-Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Sec.
-VI.2 evaluating ArtA’s overall performance.
+Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Section VI.2 evaluating ArtA’s overall performance.
 In the second way, we will only use the best-performing optimization method but expand our circuit selection to concretely answer the last research question.
@@
 During each optimization cycle, a specific termination condition (TC) is checked to reduce further computational time based on the way ArtA is used.
-Since we already obtained the ESP values for all architectures first during the third question, the highest ESP is known and can work as a TC.
+Since we first obtain the ESP values for all architectures during the third question, the highest ESP is known and can work as a TC.
 Additionally, we observed in practice that most runs reach the highest ESP sooner than 40 minutes; hence, this can be another TC.
@@
 To determine the performance of ArtA in the context of replying to the third research question, we need to define some relevant metrics.
-The optimization methods that will eventually determine ArtA’s performance should return a desired solution faster compared to searching with the brute-force approach.
+The optimization methods that will eventually determine ArtA’s performance should return a desired solution faster compared to searching with the brute-force approach of SpinQ.
 Otherwise, ArtA will be of no practical interest.
@@
 Because of that, time-to-solution is not sufficient in evaluating ArtA’s performance.
-Thus, a relative calls-to-solution metric, in Eq.
-7, is also a strong performance predictor for ArtA, especially for larger circuits.
+Thus, a relative calls-to-solution metric, in Equation 7, is also a strong performance predictor for ArtA, especially for larger circuits.
 With calls-to-solution in the numerator, we refer to the total number of times SpinQ is called to compile for each new architecture.
@@
 VII Results and Evaluation
-In this section, we will attempt to completely answer the third and fourth search questions.
+In this section, we will attempt to completely answer the third and fourth research questions.
 In both questions, we will focus on finding the architecture that obtains the highest ESP among all in the design space.
-However, before that, we briefly detail our experimental setup.
 Since all optimization methods are inherently probabilistic, multiple runs are required to create robust conclusions.
@@
 This table also provides insights into matching different method configurations (described in Appendix X.2) with specific quantum circuits.
-As explained previously, the denominator of Eq.
-6 is obtained beforehand by measuring the time taken by the brute force approach.
+As explained previously, the denominator of Equation 6 is obtained beforehand by measuring the time taken by the brute force approach.
 We observe overall that for each quantum circuit, on average, all optimization methods find the solution in less time than brute-forcing (values under 100).
@@
 Multiple optimization methods greatly outperform the random sampling (RS column), as expected.
-The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% on average of the total brute force time to find the architecture with the highest ESP.
+The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% relative time-to-solution, on average, of the total brute force time to find the architecture with the highest ESP.
 We calculated based on Table 3 that, on average, for all optimization methods, ArtA was able to decrease the exploration time by 80.6%.
@@
 More specifically, GA, with a population size of 50 and a mutation probability of 0.15, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, MA] coming close at 12.1%.
-Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results.
+Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results across the ten runs.
 To further support our findings, next, we will assess the relative calls-to-solution metric.
@@
 Since the compilation time per architecture is the same between methods, this points to a slightly (0.2% on average) longer execution time of BO [EI, MA] compared to GA [100, 0.2].
-This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Sec.
-X.2).
+This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Section X.2).
 However, the actual total runtime of ArtA will depend on the ratio between compilation and optimization time, which can change for large circuits.
@@
 Here, we use ArtA to create architectural insights into spin-qubit devices for each circuit category and suggest a universal architecture for all executed circuits.
-In Table 5, we present the values of the architectural variables, introduced in Sec.
-IV, that were found by ArtA having the highest ESP with the BO [EI, MA] optimization method alongside their ESP values.
+In Table 5, we present the values of the architectural variables, introduced in Section IV, that were found by ArtA having the highest ESP with the BO [EI, MA] optimization method alongside their ESP values.
 These circuits are organized by class and size in number of qubits within the rows of the table.
-Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Sec.
-IV.
+Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Section IV.
 The use of color accents in the table serves to enhance visual clarity, with each corresponding to the values of each variable.
@@
 Moreover, as a shuttle operation involves two quantum dots while a pulse-based operation only involves one, the shuttle-based Z rotation will have a higher chance of incurring crosstalk errors based on our crosstalk definition.
-Given these two reasons, BO has favored the shuttle-based version.
+Given these two reasons, BO has favored the pulsed-based version.
 This indicates that shuttle-based single-qubit rotations are not as sustainable as pulsed-based for large-scale circuits.
-Moving on, the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6.
+Moving on, since the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6.
 Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (BO [EI, MA]), as we can observe in the table.
@@
 For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, which has xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
-Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it is not necessary for the hardware to support parallel executions of single-qubit and two-qubit gates.
+Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it seems unnecessary for the hardware to support parallel executions of single-qubit and two-qubit gates.
 This shows that the success of real quantum circuits can come from different “angles” with reduced hardware capabilities  [^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics)..
@@
 IX Conclusion
-In this paper, we propose a comprehensive exploration of the current and future architectural design space for spin-qubit-based quantum dots quantum processors.
+In this paper, we propose a comprehensive exploration of the current and future architectural design space for quantum dot spin-qubit quantum processors.
 Our research focuses on assessing the critical architectural characteristics that could be key in ensuring high performance on such devices.
@@
 The goal of ArtA is (a) to suggest which of the seventeen optimization method configurations can find the architecture with the highest ESP the fastest compared to the brute forcing approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per circuit category.
-We have shown up to 99.1% improvement in computation times compared to the brute-force approach.
+We have shown up to 99.1% improvement in computation times compared to the brute-force approach, showing that it is possible to easily explore a vast number of designs.
 We have also provided insights into matching different optimization methods with specific quantum circuit categories.
@@
 Furthermore, our universal architecture suggestion places higher priority on parallelizing single-qubit gates rather than two-qubit gates.
+Overall, our work illustrates how the integration of DSE methodologies with optimization algorithms can be successfully utilized to offer valuable insights to designers of quantum processors.
 As for future improvements, a broader range of hyper-parameter combinations and circuit variations could enhance ArtA’s effectiveness.
-Then, with advancements in spin qubit device fabrication, new architectural variables, and values could potentially update the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA.
+Then, with advancements in spin qubit device fabrication, new architectural variables and values could potentially update the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA.
 Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world conclusions regarding developing spin qubits.

### v3

@@
 Utilizing the upgraded SpinQ compilation framework, this study explores a substantial design space comprising 29,312 spin-qubit-based architectures and applies an innovative optimization tool, ArtA (Artificial Architect), to speed up the design space traversal.
-ArtA can leverage 17 optimization configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute force approach while maintaining the same result quality.
-After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests specific and universal architectural features that provide optimal performance across the examined circuits.
-Our work demonstrates that the synergy between DSE methodologies and optimization algorithms can effectively be deployed to provide useful suggestions to quantum processor designers.
+ArtA can leverage 17 optimization configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute-force approach while maintaining the same result quality.
+After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests specific as well as universal architectural features that provide optimal performance across the examined circuits.
+Our work demonstrates that combining DSE methodologies with optimization algorithms can be effectively used to generate meaningful design insights for quantum processor development.
 I Introduction
@@
 To this day, spin qubits are at an early stage in their development, with Intel’s Tunnel Falls chip currently boasting the highest count of twelve spin qubits [34].
-Despite this, their inherent scalability advantages suggest that a robust two-dimensional design [68, 38, 8] could be scaled up relatively easily once fabrication techniques and quality improve up to a certain level.
+Despite this, their inherent scalability advantages suggest that a robust two-dimensional design [78, 41, 32, 7, 91, 38, 77] could be scaled up relatively easily once fabrication techniques and quality improve up to a certain level.
 Designing a chip, however, involves multiple architectural design choices whose impact in the future can only be fully assessed through experimental studies post-scaling the technology.
@@
 These transformations, and many more, establish SpinQ as the first compilation and DSE framework for spin qubit architectures.
-Standard DSE processes [54, 43] will be impractical time-wise, especially for such large spaces, as they rely on a brute force approach to traverse and subsequently analyze the design space.
-Our approach, however, employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods designed to automate the DSE process of SpinQ, thus taking less time than the brute force approach with the same quality of results.
-The abilities of ArtA are (a) to suggest which of the seventeen optimization configurations can find the architecture with the desired ESP the fastest compared to the brute-force approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per quantum benchmark circuit.
+Standard DSE processes [60, 47], however, will be impractical time-wise, especially for such large spaces, as they rely on a brute-force approach to traverse and subsequently analyze the design space.
+To address this, our approach employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods for automating the DSE process of SpinQ, thus taking less time than the brute-force approach with the same quality of results.
+The abilities of ArtA are (a) to suggest which of the seventeen optimization configurations can find the architecture with the desired ESP the fastest compared to the brute-force approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per quantum circuit.
 Our results demonstrate ArtA’s ability to compare all optimization techniques for each quantum circuit and obtain a solution up to 99.1% faster, on average, than brute-forcing.
@@
 Firstly, we highlight the critical need to maximize the parallelization of quantum gates rather than minimizing crosstalk between qubits.
-Secondly, we find that the communication method with shuttle operations is better in large-scale circuits than SWAPs, but for single-qubit gates, pulse-based methods rotations are preferred over shuttle-based.
-Finally, our universal architecture for all used quantum circuits suggests that the likelihood of success increases when prioritizing the parallelization of single-qubit gates instead of two-qubit gates.
-With these, we showcase that combining DSE methodologies with optimization algorithms effectively assists in forming beneficial recommendations for quantum processor designers for current and future technological stages.
-The main contributions of this paper:
+Secondly, we find that the communication method via shuttle operations is achieving higher ESP in large-scale circuits than SWAPs, but for single-qubit gates, pulse-based rotations are preferred over shuttle-based ones.
+Finally, a universal recommendation based on all tested circuits indicates that the likelihood of success increases when prioritizing the parallelization of single-qubit gates over two-qubit gates.
+This demonstrates that the combination of DSE techniques and optimization algorithms can effectively guide quantum processor designers with meaningful recommendations in the current and future stages of spin-qubit technology.
+Notably, we also observe that certain quantum circuits achieve optimal performance even with reduced hardware capabilities.
+This highlights the capability of our approach to uncover optimal architectural configurations without relying on highly complex or resource-intensive hardware designs.
+The main contributions of this paper are:
   1. 1.
@@
   2. 2.
-    The new SpinQ, the first compiler and DSE framework for spin qubit architectures.
+    The upgraded SpinQ, the first compiler and DSE framework for spin qubit architectures.
 In this version, we have updated the ESP formula to include operational, crosstalk, and decoherence-induced errors.
@@
 The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the performance of the resulting architecture.
-Then, in Section III, we motivate the need for an automated Design Space Exploration analysis for spin qubit devices and formulate four research questions for this work.
-After that, in Section IV, we define the design space consisting of various input architectural variables and establish the new DSE functionality of SpinQ.
-Another upgrade in SpinQ is the enhanced Estimated Success Probability metric, described in Section V, which is used as the Figure of Merit for our exploration.
-In Section VI, we introduce ArtA (Artificial Architect) and two new metrics to evaluate its performance.
+Then, in Section III, we motivate the need for an automated DSE analysis for spin qubit devices and formulate four research questions for this work.
+After that, in Section IV, we define the design space consisting of various input architectural variables and establish the new functionalities of SpinQ.
+Another upgrade in SpinQ is the enhanced ESP metric, described in Section V, which is used as the figure of merit for our exploration.
+In Section VI, we introduce ArtA and two new metrics to evaluate its performance.
 In the results Section VII, we first evaluate ArtA’s performance across all used quantum circuits and determine the best optimization method configuration.
-Then, we conduct a detailed analysis of the best architectural designs for each quantum circuit and form valuable insights for a universal architecture.
-We conclude our work and provide ideas for future directions in Section IX.
+Then, we conduct a detailed analysis of the best architectural designs for each quantum circuit and form valuable insights.
+We conclude our work and provide ideas for future directions in Section VIII.
 II Challenges for scaling up spin qubit devices
 Spin-qubit technologies are distinguished by their unique physical features, which position them as a highly scalable solution for quantum computing.
-The advantages of spin-qubits include a significantly smaller size — up to a thousand times less than other qubit technologies — combined with decades of semiconductor manufacturing expertise, long coherence times coupled with short gate durations, and high temperatures [76, 11, 32, 12, 79, 42, 68, 71, 77, 73].
-At the core is the quantum dot, which contains a trapped electron(s) or hole(s) to form a physical qubit [28].
+The advantages of spin-qubits include a significantly smaller size — up to a thousand times less than other qubit technologies — combined with decades of semiconductor manufacturing expertise, long coherence times coupled with short gate durations, and high operational temperatures [87, 10, 32, 11, 92, 45, 78, 81, 88, 84].
+At the core of this technology lies the quantum dot, which can contain a trapped electron(s) or hole(s) to form a physical qubit [28].
 Spin qubits are manipulated electromagnetically using multiple precision-engineered gate electrodes that facilitate either single- or two-qubit operations through exact timing of pulse sequences across various quantum dot configurations.
-Studies have expanded these systems into one-dimensional and two-dimensional arrays [32, 7], exploring different structures and material combinations.
+Studies have expanded these systems into one-dimensional and two-dimensional arrays [32, 7, 91, 38, 77], exploring different structures and material combinations.
 Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and device availability.
 Major technological hurdles are related to the so-called interconnect bottleneck [78], and various fabrication challenges towards scaling up [9, 6, 19, 18].
-On the upside, there have been significant efforts [38, 9, 68, 33, 23, 49, 52, 70] to tackle these challenges.
+On the upside, there have been significant efforts [41, 8, 78, 33, 23, 54, 58, 80, 35] to tackle these challenges.
 However, solving them can not guarantee successful quantum algorithm executions.
@@
 III Problem statement
-The need to apply design space exploration (DSE) techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2, 65, 5].
-These techniques or similar ones have been successfully applied across different levels of the quantum computing stack.
-For instance, they have been used to explore ion-trap quantum processors [45], evaluate compilation techniques [58] and qubit connectivities [41, 39, 20] for superconducting qubits.
+The need to apply DSE techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2, 74, 5].
+These techniques and similar ones have been successfully applied across different levels of the quantum computing stack which predominantly rely on brute-force exploration (i.e., exhaustive search).
+This is largely due to the inherently small design spaces in current quantum computing systems, which renders brute-force evaluation computationally feasible.
+For instance, DSE methodologies have been used to explore ion-trap quantum processors [49], evaluate compilation techniques [64] and qubit connectivities [44, 42, 20] for superconducting qubits.
 Such techniques can abstractly convert architectural characteristics into design variables and, in this way, quantize the design space, expressing current and future design possibilities that otherwise would be impossible or too time-consuming to physically realize.
-Starting from there [54, 43], the steps for a proper DSE: i) Describe the problem in terms of input variables or parameters (design choices), ii) select the performance metrics, iii) choose a global cost function as a Figure of Merit that combines different performance metrics, and iv) find models (behavioral, analytical, experimental data) that connect input parameters with metrics (or directly to the Figure of Merit).
-By using these DSE techniques, performance trends can be observed projecting current or future multidimensional design spaces, and points of optimality can be identified.
-From such an approach, it can be determined which architectural design choices (i.e., combinations of input parameter values) could provide better performance when benchmarked from small to large-scale quantum circuits.
+Starting from there [60, 47], the steps for a proper DSE: i) Describe the problem in terms of input variables or parameters (design choices), ii) select the performance metrics, iii) choose a global cost function as a figure of merit that combines different performance metrics, and iv) find models (behavioral, analytical, from experimental data) that connect input parameters with metrics or directly to the figure of merit.
+Applying these DSE techniques allows researchers to uncover performance trends, model current and future multidimensional design spaces, and identify optimal design points across a wide range of application use cases.
 Although this methodology is well established and understood in many disciplines, it can be challenging to implement in such early stages for spin-qubit technologies.
-One of the most difficult aspects is the development of a simulation framework incorporating a range of representative architectural variables adopting a large-scale perspective.
+One of the most difficult aspects is the development of a simulation framework incorporating a range of representative architectural variables with a large-scale perspective.
 Firstly, it is difficult to predict which architectural features will be relevant for many generations of devices to come, and secondly, the framework itself has to be flexible to incorporate new ones easily.
-In practice, this can be very time-consuming, and thus, typically, the design space is kept at a reasonable and, usually, small scope under specific technological assumptions.
-Considering the rapid progress of spin-qubits, the potential design space can be enlarged quickly with new considerations, rendering any previous exploration obsolete.
-Hypothesizing such a conventional DSE framework exists, it can be tedious to brute force thousands of designs and extensively analyze results.
-Hence, the relevancy of a framework automating such a DSE process, which can simplify performance evaluation of different spin-qubit architectures, is prominent.
-To capture the above problem statement, we ask the following four research questions:
+In practice, performing exhaustive exploration can be highly time-consuming; thus, even in classical systems, the design space is typically constrained to a manageable size by applying simplifying technological assumptions.
+While optimization algorithms have been successfully employed to efficiently explore significantly larger design spaces in classical computing domains [73, 12, 46, 53, 90, 65, 68, 83], such techniques remain largely unexplored in quantum computing DSE applications.
+As spin-qubit technology rapidly advances, the design space continues to grow with new considerations, making earlier explorations quickly outdated.
+Even assuming a brute-force approach already exists, the exponential growth in possible configurations can soon make it impractical to evaluate thousands of designs.
+Therefore, there is a clear need for an automated framework capable of efficiently managing this DSE process, significantly reducing the exploration time across current and future spin-qubit architectures.
+To capture the challenges of the above problem statement, we ask the following four research questions:
   1. 1.
@@
   3. 3.
-    Which optimization methods are best suited for automating the DSE per quantum cicruit? Answered in Section VII.1
+    Which optimization methods are most effective for automating the DSE process for each quantum circuit individually, as well as across all used circuits collectively? Answered in Section VII.1
   4. 4.
-    Which architectural characteristics in the selected design space are key for building high-performance spin-qubit devices for specifically given quantum circuits? Answered in Section VII.2
+    Which architectural characteristics in the selected design space are key for building high-performance spin-qubit devices for the specifically given quantum circuits? Answered in Section VII.2
 IV SpinQ and the design space
-Table 1: Summary of values of the considered architecture variables.
+Table 1: Summary of values of the considered architecture variables
 To answer the first two research questions, we conducted an analysis of relevant spin-qubit architectural properties with a long-term view and distilled them into numerous variables.
-These properties were selected to be representative of spin qubit devices as well as to create trade-offs between them.
-The latter will prevent maximizing one without causing a performance penalty through another (refer to Section V).
+These properties were selected to be representative of spin qubit devices as well as to be able to form performance-related trade-offs between them.
+The latter will prevent “maximizing” one without causing a performance penalty through another (refer to Section V).
 The defined design space is summarized in Table 1, including their possible values.
-It should be noted that the DSE analysis in this work uses either discrete or categorical variables [46], as we aim to provide performance trends and rankings between optimization methods and architectures rather than absolute comparisons.
 An architecture is then characterized by one (valid) combination of variable values.
 Below, we describe the considered design space analytically.
+It should be noted that the DSE analysis in this work uses either discrete or categorical variables [50], as we aim to provide performance trends and rankings between optimization methods and architectures.
 IV.1 Operational gate constraints
 The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types can be executed simultaneously and (b) how many of them can be parallelized at the same time step.
-To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk.
   • a)
@@
     As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel.
-For example, xy_z defines if X/Y gates can be executed simultaneously with Z gates.
+For example, xy_z defines if X or Y gates can be executed simultaneously with Z gates.
   • b)
@@
     (xyD, zD, tqgD, sD = -1, 1, 25%, 50%, 75%, 100% )
-    These variables indicate how many gates, for each type, can be parallelized, expressed as a percentage relation to the total number of qubits.
+    These variables indicate how many gates of each type can be parallelized, expressed as a percentage relation to the total number of qubits.
 For example, xyD = 25 in a 40-qubit architecture means that at most ten X or Y gates can exist in a single cycle.
 Note that the notion of a cycle refers to the basic unit of time representing one step in a sequence of gates of a quantum circuit, and each step may contain multiple gates.
-The zD and sD variables indicate the same for Z rotations and shuttle gates, respectively.
+The zD and sD variables indicate the same for Z rotations and shuttle operations, respectively.
 The tqgD variable indicates the parallelization amount for two-qubit gates.
@@
 For qubits that do not participate in a cycle, locally applied disable instructions prevent their rotation.
-Then, the Semi-Global implementation is considered with the same rotation scheme as in [30] and [51], which is essentially equivalent to the implementations of the crossbar architecture proposed of [38].
+Then, the Semi-Global implementation is considered with the same rotation scheme as in [30] and [56], which is essentially equivalent to the implementations of the crossbar architecture proposed in [41].
 In this implementation, qubit rotations are implemented semi-globally, meaning that either all qubits in odd or even column parities can be rotated at a time with the same axis and angle, and unwanted qubit rotations have to be reversed afterward by additional instructions [56].
-A Local implementation allows for any parallelized combination of gates at any angle to be executed in a cycle.
-Finally, a Sequential single-qubit gate implementation does not allow any gate parallelization, meaning that every cycle will be occupied by one gate only.
+A Local implementation enables arbitrary parallel execution of single-qubit gates, allowing any combination of rotation axes and angles to be applied simultaneously within a single cycle.
+Finally, a Sequential single-qubit gate implementation does not allow parallelizations, meaning that every cycle will be occupied by one gate only.
 It should be noted that depending on the single_qubit_impl value, only certain combinations of categories (a) and (b) are possible.
@@
 (z_rot_impl = 0, 1)
-This boolean variable also stipulates the implementation of Z rotation gates on a hardware level.
+This boolean variable stipulates the hardware-level implementation of Z rotation gates.
 The first option is a high-fidelity shuttle-based Z rotation, which is achieved with two time-sensitive qubit shuttles to and from a neighboring column [56, 30, 48, 41].
@@
 Finally, the third adds the second diagonal in each square (i.e., each qubit is coupled with eight neighboring qubits and with three at the corners, degree = 8).
-Figure 1, presents the coupling graphs considered with color-coded edges for each of the three types.
-We have now reviewed all the variables related purely to hardware.
-Next, we will consider two variables associated also with the compilation process.
-More specifically, they change the routing process but based on the hardware communication constraints.
-Figure 1: The different coupling graphs considered in the design space.
-The coupling graph consists of the black line edges when degree = 4, black and green edges when degree = 6, and all edges when degree = 8.
-IV.4 Routing methods:
+Figure 1(a) presents the coupling graphs considered with color-coded edges for each of the three types.
+The actual dimensions of each coupling graph, and thus the overall size of the quantum architecture, are determined by the number of qubits required by the quantum algorithm under execution.
+Figure 1: (a) The different coupling graphs considered in the design space.
+The coupling graph consists of the black line edges when degree = 4, black and green edges when degree = 6, and all edges when degree = 8 (b) Schematic overview of the crossbar architecture, taken from [56], with the various shared operational control lines: vertical (column line, CL), horizontal (row line, RL), and diagonal (qubit line, QL).
+These lines are used with precise pulse sequences and shared among multiple sites, sixteen quantum dots in this figure, to perform operations on qubits [41, 30, 48] under specific operational constants.
+Here, the qubits (green circles with numbers) are initialized in a checkerboard pattern.
+(c) Abstraction of the crossbar architecture taken from [55] representing the coupling graph between qubits with degree = 4.
+Each circle represents a quantum dot, and each edge represents a coupling link signifying allowed interactions.
+To further ease the understanding of the architectural landscape of the design space, in Figure 1(b), we depict the crossbar architecture [41] as a real-world example of an architecture represented in the design space.
+In this figure, the operational lines confine sixteen quantum dots, eight of which are occupied by spin qubits.
+The operational gate constraints in this architecture arise from the shared-control lines and other physical properties as described in [41, 30, 48, 56].
+Among these constraints, the coupling connectivity is particularly relevant, as it corresponds to a square grid with degree = 4 in this implementation.
+Consequently, the topology of the crossbar architecture can be abstracted as a two-dimensional grid, as depicted in Figure 1(c).
+We have now reviewed all variables that pertain solely to the hardware.
+Next, we examine two additional variables that influence the compilation process.
+Specifically, these variables can alter the routing methodology based on the underlying hardware communication constraints.
+IV.4 Routing methods via shuttling:
 (router = 0 $\rightarrow$ Shuttle-based SWAP, 1 $\rightarrow$ beSnake)
-There are two routing algorithm choices.
+There are two available algorithmic options.
 The first one, introduced in [56, 48], is a shuttle-based SWAP routing algorithm.
 In particular, this algorithm was tailored around the unique constraints of the crossbar architecture [41], which necessitated the maintenance of the checkerboard physical qubit pattern to achieve a time-efficient compilation process.
-One of the advantages of this algorithm is, in fact, the maintenance of the checkerboard pattern, which in turn keeps the crosstalk effect low and enables the Semi-Global rotations scheme exclusively.
+One of the advantages of this algorithm is, in fact, the maintenance of the checkerboard pattern, which in turn keeps the crosstalk interference low and exclusively enables the Semi-Global rotations scheme.
 The second routing option is called beSnake [55], and it is capable of freely shuttling qubits around any topology and in any direction, handling complex routing scenarios involving parallelized gates.
@@
 (SWAP_opt = 0, 1)
-This binary variable concerns an optional functionality of the beSnake routing algorithm, which can replace a sequence of shuttles with a SWAP operation under certain conditions [50].
+This binary variable concerns an optional functionality of the beSnake routing algorithm, which can replace a sequence of shuttles with a SWAP gate under certain conditions [55].
 These conditions are satisfied when the accumulated fidelity of a parallelized shuttle sequence exceeds the fidelity of only executing a SWAP on that particular location.
@@
 All these variable values act as input into the upgraded SpinQ housing its new configurable compiler, making it the first compilation and DSE framework for spin-qubit architectures.
-After this upgrade, we are able to provide SpinQ with multiple combinations of variable values, compile each input circuit iteratively for all valid architectures, and calculate and store the same performance metrics used in [51].
-These are gate overhead, circuit depth overhead, and Estimated Success Probability (ESP) of the compiled quantum circuit(s).
+After this upgrade, we are able to provide SpinQ with multiple combinations of variable values, compile each input circuit iteratively for all valid architectures, and calculate and store multiple performance metrics, such as the ones used in [56].
+These are gate overhead, circuit depth overhead, and ESP of the compiled quantum circuit(s).
 As mentioned, not all combinations of variable assignments are valid, as there are interdependencies between them; hence, SpinQ automatically filters the design space to allow only valid combinations.
@@
 This tool is crucial, given the absence of real devices for testing.
-The verification functions have been enhanced to handle all valid architectures via our two-step method, which thoroughly scrutinizes circuits at each compilation stage [51].
+For this work, the verification functions have been enhanced to handle all valid architectures via our two-step method [56], which thoroughly scrutinizes circuits at each compilation stage.
 V Figure of merit
-To estimate the performance of any quantum processor, a prevalent method when the device is not physically available is through the Estimated Success Probability (ESP) [57, 47] extracted from the compiled circuit.
-This approach is by far the most efficient compared to Schrodinger-Feynman, tensor networks, or noise model simulations as all become exponentially slower with larger circuits and/or higher qubit connections on the device [51].
-The choice of the figure of merit is a crucial step to ensure fair results, as every design variable needs to be directly linked to the resulting value, along with their trade-offs.
-This will prevent ArtA (introduced later in Section VI) from maximizing some variable(s) and, by doing so, unfairly favoring specific architectures.
-For example, it will not be fair to be able to maximize the parallelization degree of all gates without causing some negative impact due to increased crosstalk.
-Such a consideration also expresses more accurately design decisions in real experimental quantum processors.
-In this work, we start from the ESP definition of [51] given in Equation 1 and expand it to reflect all previous considerations, given in Equation 2, with now three components in total: a gate operations term, a crosstalk term, and a decoherence term.
+The selection of an appropriate figure of merit is a critical step to ensure fair and objective evaluation of quantum processor architectures.
+It must reflect how specific design features influence circuit performance, allowing each architecture to be characterized based solely on its inherent trade-offs and capabilities.
+To achieve this, each design variable must contribute directly to the final metric, such that the interplay between performance gains and associated costs (e.g., crosstalk or decoherence) is properly captured.
+This prevents the optimization tool ArtA (introduced in Section VI) from disproportionately favoring certain architectures by, for instance, maximizing the degree of gate parallelization without accounting for the negative effects it may introduce, such as increased crosstalk.
+To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk.
+Such considerations also expresses more accurately design decisions in real experimental quantum processors.
+When a quantum processor is not physically available, a common method for estimating its performance is to compute the Estimated Success Probability (ESP) [63, 51, 57], derived from the compiled quantum circuit.
+This method is significantly more computationally efficient than alternatives such as Schrödinger-Feynman simulation, tensor network contraction, or noise model simulations, all of which scale poorly with increasing circuit size and qubit connectivity [56, 57].
+In this work, we build upon the ESP definition introduced in [56], shown in Equation 1, and extend it to incorporate all previously discussed architectural considerations.
+The revised figure of merit, presented in Equation 2, now consists of three components: a gate operations term, a crosstalk term, and a decoherence term.
 $ESP=\prod_{k}\prod_{i}{F_{i,k}}$ (1)
-where $k$ represents the $k$th time step and $i$ the $i$th gate in the $k$th time step.
-$ESP=(\prod_{k}\prod_{i}{F_{i,k}})\cdot(\prod_{k}\prod_{i,j}C_{i,j,k})\cdot(1-e^{(-t/T_{2})})^{N}$ (2)
-where $k$ represents the $k$th time step, $i$ the $i$th gate in the $k$th time step and $j$ represents the $j$th gate that inflicts crosstalk with the $i$th gate.
-$F$ is the fidelity and $C$ the crosstalk of the corresponding gates.
-The last term, adapted from [30], introduces the decoherence-induced errors, which represent the probability of all qubits staying coherent during the execution of the circuit.
-$T_{2}$ is the decoherence time, $t$ the duration time of the circuit, and $N$ the total number of qubits.
+where $k$ represents the $k$th time step and $i$ the $i$th gate in the $k$th time step. $F$ is the fidelity of the corresponding gates.
+$ESP=(\prod_{k}\prod_{i}{F_{i,k}})\cdot(\prod_{k}\prod_{i,j}C_{i,j,k})\cdot{\color[rgb]{0,0,0}(e^{(-t/T_{2}^{*})})^{N}}$ (2)
+where $C$ denotes the crosstalk between the $i$th and $j$th gates during the $k$th time step.
+The last term, adapted from [30], introduces the decoherence-induced errors, which represent the probability of all qubits staying coherent during the execution of the circuit’s execution.
+$T_{2}^{*}$ is the decoherence time, $t$ the total duration time of the circuit, and $N$ the total number of qubits.
 We will discuss each of the terms in detail in the next three subsections.
 V.1 Gate operations term
-The choice of typical operational fidelities of each gate, coined by the $F_{i,k}$ term, plays a large role in the probability of success of quantum circuits.
-In our analysis, we consider these fidelities to be constants within the design space rather than variables.
-This is because determining the optimal fidelity value for enhancing the ESP is straightforward; higher operational fidelity invariably leads to improved performance.
+The selection of typical operational fidelities for each gate, denoted by the $F_{i,k}$ term, plays a significant role in determining the success probability of quantum circuits.
+In our analysis, we consider these fidelities to be constants within the design space, rather than variables.
+This is because determining the optimal fidelity value for increasing the figure of merit (i.e., ESP) is straightforward; higher operational fidelity invariably leads to improved performance.
 Since we are conducting an exploration, we take a highly optimistic approach to the state-of-the-art [59, 86, 52] values for each gate type [^1] Another reason we do so is to avoid numeric underflows where there is a loss of accuracy in numerical calculations if ESP becomes smaller than the smallest positive representable value in a programming language’s floating-point arithmetic.:
@@
 V.2 Crosstalk term
-Crosstalk has been sparsely studied for spin qubits.
-This is because the crosstalk effects when executing multiple gates in parallel in close proximity are still very complex to model accurately [66, 29].
-Nonetheless, the presence of crosstalk is expected to have a large influence on quantum computer performance in the future, and current proposals are already suggesting techniques to mitigate it in simple forms [66].
+Crosstalk in spin qubit systems has received limited attention in the literature, primarily due to the complexity of accurately modeling the effects that arise when multiple gates are executed simultaneously in close physical proximity [75, 29, 37].
+Nonetheless, the presence of crosstalk is expected to have a large influence on quantum computer performance in the future, and current proposals are already suggesting techniques to mitigate it in simple occurrences [75].
 For our purposes, crosstalk creates necessary trade-offs between architectural variables, and, as explained before, this is essential for a fair ESP calculation across architectures.
@@
 where $n$ is the number of the total direct links (i.e., nearest-neighbors) on the topology between the $i$th and the $j$th operation. Similarly to before, $F$ corresponds to the operational fidelity.
-A schematic overview of crosstalk effects is given in Figure 3.
-We see two two-qubit gates acting within the red ovals between Q-1 and Q-2, and Q-3 and Q-4, respectively.
-In this hypothetical square grid with one diagonal topology, crosstalk occurs in all three edges marked with a flash, as they directly connect the qubits of these two operations.
-Figure 3: Schematic overview of cross-talk when executing two-qubit gates.
-Dashed black lines indicate a direct connection between two qubits, the red ovals indicate an ongoing two-qubit gate, and the blue flash icons indicate contributing crosstalk effects.
+A schematic overview of crosstalk effects is presented in Figure 3.
+The diagram illustrates two two-qubit gates, represented by red ovals, acting between qubits Q-1 and Q-2, and Q-3 and Q-4, respectively.
+In this hypothetical architecture with degree = 4 connectivity, crosstalk arises along the three edges marked with a lightning symbol, as the involved qubits are directly connected to one another.
+Figure 3: Schematic overview of crosstalk during the execution of two two-qubit gates.
+Dashed black lines represent direct qubit connections, red ovals indicate active two-qubit gates, and blue lightning icons denote edges where crosstalk effects occur.
 V.3 Decoherence term
-The last term of Equation 2 represents the decoherence of quantum information, which creates a strong trade-off between parallelization and crosstalk effects.
-Since $N$ and $T_{2}$ are given, we need to define the duration time of the circuit $t$:
+The last term of Equation 2 represents the decoherence of quantum information, which creates a strong architectural trade-off between parallelization and crosstalk effects.
+Since $N$ and $T_{2}^{*}$ are given, we need to define the duration time of the circuit $t$:
 $t=\sum_{k=1}^{M}\text{max}(D_{i,k})$ (5)
@@
 The motivation behind this selection of durations, once again, is for purposes of architectural comparisons.
-We took reference for our two-qubit gate (i.e., $\sqrt{SWAP}$) from [48, 75], for the $Shuttle$ from [30] which is half the shuttle-based Z rotations [51], for the pulsed-based single-qubit gates from [53], and the $T_{2}$ from [31].
+We took reference for our two-qubit gate (i.e., $\sqrt{SWAP}$) from [52, 86], for the $shuttle$ from [30] which is half the shuttle-based Z rotations [56], for the pulse-based single-qubit gates from [59], and for $T_{2}^{*}$ from [31].
 Table 2: Overview of all gate duration and the decoherence time used.
-VI ArtA (Artificial Architect)
-Figure 4: Schematic overview of the initialization, optimization, and output cycle when using ArtA (red box) together with the upgraded SpinQ framework (blue box).
-From the aforementioned transformation of the SpinQ compilation framework to a Design Space Exploration (DSE) framework, we are able to brute force through all valid combinations of the input variable values.
+VI ArtAArtificial Architect
+From the aforementioned transformation of the SpinQ compilation framework to a fully functional DSE framework, we are able to brute-force through all valid combinations of the input variable values.
 However, as we discussed, this can be a time-consuming process, especially when the defined design space is large.
 Therefore, a specialized tool is needed to automate the DSE process by traversing the space in a clever and faster way.
-Ultimately, this will result in fewer designs being tried until desired ESP values or other metrics are reached.
-To do that, we introduce ArtA, a tool built into the SpinQ framework containing multiple optimizers.
-Its goal is to give answers to (a) which optimization technique performs best on the task of finding an architecture that achieves the desired ESP and (b) which architecture provides the highest ESP, both per quantum circuit and universally.
+As a result, fewer design iterations will be tried to achieve the same ESP (range) or other specified performance metrics.
+Figure 4: Schematic overview of the initialization, optimization, and output cycle when using ArtA (red box) together with the upgraded SpinQ framework (blue box).
+To do that, we introduce ArtA, a tool integrated into the SpinQ framework that incorporates multiple optimization techniques.
+Its purpose is twofold: (a) to determine which optimization method is most effective at identifying architectures that achieve the desired ESP for each quantum circuit, and (b) to identify the architecture that yields the highest ESP both on a per-circuit basis and across all circuits collectively (i.e., a universal architecture).
 The abstracted ArtA process is depicted in Figure 4.
 The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Section IV, the optimization method(s) to traverse the space, and an (optional) termination condition.
-Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture (in the form of input variable values) from the available “population” of architectures – a subset of the complete design space determined by the optimization algorithm itself.
-This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired value of a specific metric is reached (ESP in this work).
-Upon termination, the resulting architecture, its ESP, and ArtA metrics are stored.
-In the following sections, we will describe the TC and ArtA metrics in more detail.
+Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture in the form of input variable values from the available “population” of architectures – a subset of design space selected by the optimization algorithm itself.
+This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired range/value of a specific metric is reached (ESP in this work).
+Upon termination, the resulting architecture, its ESP calculation, and ArtA metrics are stored.
+In the following sections, we will describe the TCs and ArtA metrics in more detail.
 In the Appendix X.2, we analytically discuss the five optimization methods used and their hyper-parameter variations, totaling seventeen different optimization configurations.
@@
 To answer the last two research questions, we will use ArtA’s (a) and (b) capabilities.
-In the first one, we answer the third research question by testing all the implemented optimization methods to determine which finds the highest ESP the fastest for each quantum circuit.
-Therefore, we will compile each circuit for all $29,312$ architectures and store their ESP.
-Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Section VI.2 evaluating ArtA’s overall performance.
-In the second way, we will only use the best-performing optimization method but expand our circuit selection to concretely answer the last research question.
-VI.1 Termination Condition (TC)
-During each optimization cycle, a specific termination condition (TC) is checked to reduce further computational time based on the way ArtA is used.
-Since we first obtain the ESP values for all architectures during the third question, the highest ESP is known and can work as a TC.
+With the first one, we will answer the third research question, in Section VII.1, by testing all the implemented optimization methods to determine which finds the highest ESP the fastest for each quantum circuit.
+Therefore, first we need to compile each circuit for all $29,312$ architectures and store their ESP for later use.
+Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Section VI.2, evaluating ArtA’s overall performance.
+With the second capability of ArtA, we will only use the best-performing optimization method, but expand our circuit selection to concretely answer the last research question, in Section VII.2.
+VI.1 Termination condition
+During each optimization cycle, a specific TC is checked to reduce further computational time based on which way, (a) or (b), ArtA is used.
+Since we first obtain the ESP values for all architectures during the third question, as mentioned before, the highest ESP is known and can work as a TC.
 Additionally, we observed in practice that most runs reach the highest ESP sooner than 40 minutes; hence, this can be another TC.
@@
   • If the method’s run time exceeds 40 minutes.
-For the second use of ArtA, which answers the fourth research question, we investigate the returned architectural designs by using the chosen optimization method derived from answering the third question.
-We find in practice that the corresponding chosen optimization algorithm did not exceed trying 23.4% of the total $29,312$ architectures in the worst case before finding the one with the highest ESP.
+For the second use of ArtA, which answers the last research question, we investigate the returned architectural designs by using the best optimization method derived from answering the third question.
+As we will see later in Section VII.1, we find that the corresponding chosen optimization algorithm did not exceed 23.4% of the total $29,312$ architectures in the worst case before finding the one with the highest ESP.
 Therefore, when answering the last research question, the TC triggers:
@@
 VI.2 Metrics evaluating ArtA
-To determine the performance of ArtA in the context of replying to the third research question, we need to define some relevant metrics.
-The optimization methods that will eventually determine ArtA’s performance should return a desired solution faster compared to searching with the brute-force approach of SpinQ.
+To determine the performance of ArtA, in the context of replying to the third research question, we need to define some relevant metrics.
+The optimization methods that will eventually determine ArtA’s performance should return a solution faster compared to searching with the native brute-force approach of SpinQ.
 Otherwise, ArtA will be of no practical interest.
 We thus use the relative time-to-solution as a metric in Equation 6.
-The numerator includes the total runtime of both the optimization method and compilation until the desired result is reached (or the TC is triggered), and the denominator refers to the total runtime to brute force the entire design space.
+The numerator includes the total runtime of both the optimization method and compilation until the desired result is reached (or the TC is triggered), and the denominator refers to the total runtime to brute-force the entire design space.
 […]
-When the average compilation time increases due to larger quantum circuits, the execution time of the optimization method slowly becomes less important within the total runtime.
-Because of that, time-to-solution is not sufficient in evaluating ArtA’s performance.
-Thus, a relative calls-to-solution metric, in Equation 7, is also a strong performance predictor for ArtA, especially for larger circuits.
+When the average compilation time increases due to larger quantum circuits, the execution time of the optimization method becomes relatively less significant in the total runtime.
+This happens because, as circuits grow in size, the compilation time per architecture dominates the total runtime, while the time taken by the optimization method remains approximately constant regardless of circuit size.
+Therefore, time-to-solution is not sufficient in evaluating ArtA’s performance, as it becomes increasingly influenced by compilation time, rather than the efficiency of the optimization method itself.
+To address this, we also introduce a relative calls-to-solution metric in Equation 7.
 With calls-to-solution in the numerator, we refer to the total number of times SpinQ is called to compile for each new architecture.
-In case the optimization method attempts the same architecture twice or more, the compilation will be skipped since the results are already stored from the first call.
-Finally, the denominator represents the brute force approach and is the total number of architectures in the design space, which essentially equals $29,312$.
+In case the optimization method attempts the same architecture twice or more, the circuit compilation will be skipped since the results are already stored from the first call.
+Finally, the denominator represents the brute-force approach and is the total number of architectures in the design space, which essentially equals $29,312$.
 […]
@@
 In this section, we will attempt to completely answer the third and fourth research questions.
-In both questions, we will focus on finding the architecture that obtains the highest ESP among all in the design space.
+In both questions, we will focus on finding the architecture that obtains the highest ESP among all within the design space.
 Since all optimization methods are inherently probabilistic, multiple runs are required to create robust conclusions.
-Due to limited computational resources [^2] All time metrics were retrieved by running ArtA with Python 3.10.8 single-threaded on a 2017 MacBook Pro, using a 3.1 GHz Quad Core Intel i7 processor and 16 GB of RAM., we ran each method for each hyper-parameter combination ten times alongside the TC.
+Due to limited computational resources [^2] All time metrics were retrieved by running ArtA with Python 3.10.8 single-threaded on a 2017 MacBook Pro, using a 3.1 GHz Quad Core Intel i7 processor and 16 GB of RAM., we ran each method for each hyper-parameter combination ten times alongside the TCs.
 Besides the mean performance, we will also report the worst performing results for reference.
@@
 A minus symbolizes that there was at least one run (out of the ten) in which the top ESP value was not reached in time (TC = 40 minutes).
-These results are colored red, as well as runs where the time taken exceeds the brute force time (values over 100).
+These results are colored red, as well as runs where the time taken exceeds the brute-force time (values over 100).
 The green cells indicate the top five configurations of optimization methods for a given quantum circuit.
@@
 The results of the relative time-to-solution metric for finding the architecture with the highest ESP are given in Table 3.
-This table also provides insights into matching different method configurations (described in Appendix X.2) with specific quantum circuits.
-As explained previously, the denominator of Equation 6 is obtained beforehand by measuring the time taken by the brute force approach.
+This table also provides insights into matching different optimization method configurations (described in Appendix X.2) with specific quantum circuits.
 We observe overall that for each quantum circuit, on average, all optimization methods find the solution in less time than brute-forcing (values under 100).
-For a better interpretation of the numbers on the table, a 50 means the particular optimization configuration of ArtA takes half the time of brute forcing all architecture, and 100 means that it takes the same time.
+For a better interpretation of the numbers on the table, a 50 means the particular optimization configuration of ArtA takes half the time of brute-forcing all architecture, and 100 means that it takes the same time.
 Multiple optimization methods greatly outperform the random sampling (RS column), as expected.
-The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% relative time-to-solution, on average, of the total brute force time to find the architecture with the highest ESP.
+The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% relative time-to-solution, on average, of the total brute-force time to find the architecture with the highest ESP.
 We calculated based on Table 3 that, on average, for all optimization methods, ArtA was able to decrease the exploration time by 80.6%.
@@
 In a performance order, we can observe GA being first, followed by BO, ACO, and, lastly, SA.
-More specifically, GA, with a population size of 50 and a mutation probability of 0.15, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, MA] coming close at 12.1%.
+More specifically, GA, with a population size of 100 and a mutation probability of 0.2, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, Ma] coming close at 12.1%.
 Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results across the ten runs.
@@
 We calculated based on Table 4 that, on average, ArtA’s optimization method configurations can decrease the compiler calls by 79.9%.
-We can observe, once again, that BO, followed by GA, obtained the best performance compared to other optimization methods.
+We can observe, once more, that BO, followed by GA, obtained the best performance compared to other optimization methods.
 Lastly, in 91.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, showing results consistency again.
@@
 Since the compilation time per architecture is the same between methods, this points to a slightly (0.2% on average) longer execution time of BO [EI, Ma] compared to GA [100, 0.2].
-This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Section X.2).
+This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Appendix X.2).
 However, the actual total runtime of ArtA will depend on the ratio between compilation and optimization time, which can change for large circuits.
@@
 VII.2 Architecture analysis
-Table 5: Architectures returned by ArtA with BO [EI, MA] optimization method for a range of quantum circuits.
+Table 5: Architectures returned by BO [EI, Ma] optimization method for a range of quantum circuits (refer to Appendix X.1).
 Results are color-coded, with color schemes following the categories introduced in Section IV.
+Final-row averages omit -1 values.
 We move on to address the last research question after having selected the best-performing optimization method in the previous section.
 Here, we use ArtA to create architectural insights into spin-qubit devices for each circuit category and suggest a universal architecture for all executed circuits.
-In Table 5, we present the values of the architectural variables, introduced in Section IV, that were found by ArtA having the highest ESP with the BO [EI, MA] optimization method alongside their ESP values.
-These circuits are organized by class and size in number of qubits within the rows of the table.
+In Table 5, we present the values of the architectural variables, introduced in Section IV, that were found by ArtA having the highest ESP with the BO [EI, Ma] optimization method alongside their respected ESP values.
+These circuits are organized by class and number of qubits within the rows of the table.
 Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Section IV.
@@
 Moreover, as a shuttle operation involves two quantum dots while a pulse-based operation only involves one, the shuttle-based Z rotation will have a higher chance of incurring crosstalk errors based on our crosstalk definition.
-Given these two reasons, BO has favored the pulsed-based version.
-This indicates that shuttle-based single-qubit rotations are not as sustainable as pulsed-based for large-scale circuits.
+Given these two reasons, BO has favored the pulse-based version.
+This indicates that shuttle-based single-qubit rotations are not sustainable compared to pulse-based ones for large-scale circuits as they progressively contribute more noise.
 Moving on, since the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6.
-Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (BO [EI, MA]), as we can observe in the table.
-Additionally, xy_z_tqg can also be conceptually equivalent to xy_tqg and z_tqg since the pulse-based Z rotation is selected.
+Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (i.e., BO [EI, Ma] in this case), as we can observe in the table.
+Additionally, xy_z_tqg can also be conceptually equivalent to xy_tqg and z_tqg since pulse-based Z rotations are selected.
 Therefore, whenever xy_z_tqg is 1 or xy_tqg and z_tqg is 1, it means the architecture(s) which allow(s) parallelization of all single- and two-qubit gates obtains the highest ESP.
@@
 However, since such exploration freedom was given in the design space, the reasons for the mismatches could be explained by the internal characteristics of the circuits themselves.
-For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, which has xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
-Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it seems unnecessary for the hardware to support parallel executions of single-qubit and two-qubit gates.
-This shows that the success of real quantum circuits can come from different “angles” with reduced hardware capabilities  [^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics)..
+For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, as observed from their gate parallelization constraints xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
+Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it appears that hardware support for parallel execution of single- and two-qubit gates is not essential.
+This highlights that the success of real quantum circuits can stem from different architectural “angles,” even under reduced hardware capabilities [57].
+This exemplifies the strength of our approach: it enables the identification of optimal architecture configurations without requiring the most complex or fabrication-intensive hardware designs.
+[^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics).
 Another strong preference appears in the beSnake router, observed by the dark green 1 values in the router column.
@@
 In fact, the decoherence-induced errors increase more severely (exponentially) with time, based on Equation 4, than crosstalk, especially in circuits with high qubit and gate counts, which explains the prevailing preference for maximizing parallelization.
-The same can be observed in our universal architecture recommendation in the last row, which performs best on average for all the circuits examined.
+The same can be observed in our universal architecture recommendation in the last row when rounding the numbers, which performs best on average for all the circuits examined.
 Additionally, the operational fidelity, which is directly linked to the gate overhead of compilation, can also severely impact the final ESP for circuits with high gate counts, which in turn explains the need to add the least additional gates possible during compilation through the use of beSnake [55] and local single-qubit implementation.
-Having said that, Bernstein Vazirani, QV8, and QV10 circuits preferred the global implementation, as these circuits consist of a significant number of similar single-qubit gates being executed at the same time step.
+Having said that, Bernstein Vazirani, QV8, and QV10 circuits preferred the global implementation, as these circuits consist of a significant number of single-qubit gates (with the same angles) being executed at the same time step.
 Weak preference
 Overall, we can observe that the single-qubit operations and shuttling prefer a higher degree of parallelization than the two-qubit gates.
-Considering that xyD and zD have the same meaning, their combined averages (excluding -1 values) from the last row of Table 5 is 72.7% and the sD to 66.7%.
-This compares to an average of 52.6% for the two-qubit gates.
-Therefore, prioritizing the simultaneous operation of single-qubit gates (and shuttling) over two-qubit gates is more important in spin-qubit architectures.
+Considering that xyD and zD have the same meaning due to zri = 1, their combined averages from the last row of Table 5 is 72.7%.
+This compares to an average of 52.6% for the two-qubit gates and 66.7% for sD.
+Therefore, prioritizing the simultaneous operation of single-qubit gates (and shuttling) over two-qubit gates is more important.
 The swap_opt variable seems to show a preference for supporting SWAP gates for lower qubit counts.
@@
 As we explained before, the effects of having longer circuit execution times become more severe with larger circuits; hence, the extra SWAPs can outweigh the gained fidelity they initially offer.
-Combining this conclusion for large-scale circuits and the strong preferences for shuttling freedom (beSnake router) indicates that the shuttle operation is a key communication method for spin-qubit devices.
-VIII Acknownledgement
-This work is part of the research program OTP with project number 16278, which is (partly) financed by the Netherlands Organisation for Scientific Research (NWO).
-CGA acknowledges funding from the Spanish Ministry of Science, Innovation and Universities through the Beatriz Galindo program 2020 (BG20-00023) and from the European ERDF under grant PID2021-123627OB-C51.
-IX Conclusion
+Combining this conclusion for large-scale circuits with the observed strong preference for shuttling flexibility, enabled by the beSnake router, suggests that shuttle operations constitute a key communication mechanism in spin-qubit architectures for the future.
+VIII Conclusion
 In this paper, we propose a comprehensive exploration of the current and future architectural design space for quantum dot spin-qubit quantum processors.
 Our research focuses on assessing the critical architectural characteristics that could be key in ensuring high performance on such devices.
-To this end, we present an upgraded version of the SpinQ compilation framework for spin qubit architectures in which architectural variables can configure internal compilation passes, making it the first Design Space Exploration (DSE) framework.
+To this end, we present an upgraded version of the SpinQ compilation framework [56] for spin qubit architectures in which architectural variables can configure internal compilation passes, resulting in the first Design Space Exploration (DSE) framework.
 After defining our design space, consisting of $29,312$ architectures, we propose a multi-optimization-based tool, ArtA (Artificial Architect), to automate the DSE process.
-The goal of ArtA is (a) to suggest which of the seventeen optimization method configurations can find the architecture with the highest ESP the fastest compared to the brute forcing approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per circuit category.
-We have shown up to 99.1% improvement in computation times compared to the brute-force approach, showing that it is possible to easily explore a vast number of designs.
+The goal of ArtA is twofold: (a) to identify which of the seventeen optimization configurations can most efficiently discover the architecture that achieves the highest ESP, compared to brute-force exploration; and (b) to determine the architectural design characteristics that are critical for constructing high-performance spin-qubit devices across different circuit categories.
+We have shown up to 99.1% improvement in computation times compared to brute-forcing, showing that it is possible to easily explore a vast number of designs.
 We have also provided insights into matching different optimization methods with specific quantum circuit categories.
-Bayesian optimization with an Expected Improvement acquisition function and a Matérn kernel was chosen as the best optimization method due to its effectiveness in managing higher qubit counts.
+It should be noted, that the underlying goal of our study was not to exhaustively benchmark all possible optimization methods, but rather to introduce the concept of applying design space exploration to quantum computing systems – with a particular focus on spin-qubit architectures – by employing a diverse and representative selection of methods from different optimization families.
+Bayesian optimization with an Expected Improvement acquisition function and a Matérn kernel was chosen as the best method due to its effectiveness in managing higher qubit counts.
 After that, we proceed to automate the DSE process for 44 quantum circuits of up to 258 qubits within the 29,312-architecture design space.
@@
 Lastly, the shuttle operation is preferred over costly SWAP operations for communication purposes, especially for large-scale circuits.
-Contrary to that, shuttle-based single-qubit gates such as Z rotations are not preferred over pulsed-based.
+Contrary to that, shuttle-based single-qubit gates such as Z rotations are not preferred over pulse-based.
 Furthermore, our universal architecture suggestion places higher priority on parallelizing single-qubit gates rather than two-qubit gates.
-Overall, our work illustrates how the integration of DSE methodologies with optimization algorithms can be successfully utilized to offer valuable insights to designers of quantum processors.
+Overall, our work demonstrates how the integration of DSE methodologies with optimization algorithms can uncover valuable architectural insights that would be difficult, if not impossible, to foresee while suggesting points of optimality with reduced hardware complexity.
+However, it is important to acknowledge that while these findings highlight promising architectural directions, several of these design choices – such as high degrees of parallelization, connectivity, and flexible shuttling strategies – currently pose significant engineering challenges.
+Implementing fully local control, for instance, demands high fabrication uniformity and precise calibration, which is non-trivial at larger scales.
+Likewise, extensive shuttling flexibility requires precise individual gate addressing, which is experimentally demanding.
+Nonetheless, the purpose of this work is to chart a clear path forward, identifying architectural features that, while ambitious, can guide future efforts toward scalable and performant spin-qubit processors.
 As for future improvements, a broader range of hyper-parameter combinations and circuit variations could enhance ArtA’s effectiveness.
-Then, with advancements in spin qubit device fabrication, new architectural variables and values could potentially update the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA.
-Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world conclusions regarding developing spin qubits.
-This would allow for operational fidelities to become variables as well and, consequently, add a second dimension to the Figure of Merit (ESP).
+Then, with advancements in spin qubit device fabrication, new architectural possibilities could potentially expand the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA.
+Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world predictions regarding developing spin qubits.
+For instance, this would allow for operational fidelities to become variables.
 Also, some of the assumptions made in this work could be fine-tuned as the field progresses.
 It is crucial to acknowledge the impact of these assumptions on the observations presented.
-For instance, we assumed that the fidelity of pulse-based Z rotations is equivalent to that of a single shuttle operation.
+For instance, we assumed the fidelity of pulse-based Z rotations to be equivalent to that of a single shuttle operation.
 This assumption should not be construed as either favorable or unfavorable; rather, it is a foundational premise upon which ArtA bases its architectural “recommendations”.
 Should empirical evidence later demonstrate that shuttle operations exhibit significantly higher fidelity compared to pulse-based Z rotations or the crosstalk model depends on more factors, it may necessitate a reevaluation of preferences.
-Further analysis of architectural patterns with similar quantum circuit characteristics can reveal more trends.
-For instance, quantum circuits with high two-qubit gate percentage ranges can perform best when the qubit connectivity of the architecture is above a certain degree.
-These kinds of statements or more can be made for a plethora of circuit characteristics [3, 51] and, in doing so, create a general matching between algorithm categories and architectural designs.
-Once this is done, it is possible to make the reversed suggestion of which architecture will perform best for an input circuit of interest once categorized without needing to repeat the DSE process.
+Further analysis of architectural patterns associated with specific quantum circuit characteristics may reveal additional trends and enable systematic categorization.
+For example, circuits with a high proportion of two-qubit gates (e.g., 90% or more) may exhibit optimal performance only when the qubit connectivity exceeds a certain threshold.
+Insights like these, based on various circuit properties [3, 56], can inform the construction of a generalized lookup table that maps algorithmic profiles to architectural configurations.
+Such a resource would enable rapid prediction of high-performing architectures for new circuits, eliminating the need to re-run the full DSE process.
+IX Acknownledgement
+This work is part of the research program OTP with project number 16278, which is (partly) financed by the Netherlands Organisation for Scientific Research (NWO).
+CGA acknowledges funding from the Spanish Ministry of Science, Innovation and Universities through the Beatriz Galindo program 2020 (BG20-00023) and from the European ERDF under grant PID2021-123627OB-C51.
+Declarations
+There are no competing interests between the authors.